### v4

# Nonlinear Superconducting Magnetoelectric Effect

Abstract
A supercurrent flow can induce a nonvanishing spin magnetization in noncentrosymmetric superconductors with spin-orbit interaction. Often known as the non-dissipative magnetoelectric effect, these are most commonly found at linear order in supercurrent flow. Here, we argue that a nonlinear superconducting magnetoelectric effect (NSM) can naturally manifest in altermagnet/superconductor (ALM/SC) heterostructures: NSM manifests as a spin polarization generated as a second-order response to a driving supercurrent. Strikingly, we find NSM is the leading order magnetization response in ALM/SC heterostructures and survives even in the presence of centrosymmetry; $C_{4}\mathcal{T}$ symmetry in altermagnets zeroes both the equilibrium magnetization as well as out-of-plane linear magnetoelectric response. This renders NSM a powerful electric and non-dissipative means of controlling magnetization in ALM/SC heterostructures, a promising platform for superconducting spintronics.
Introduction.—Unlike the normal metallic state, the intrinsically non-dissipative flow of current in superconductors enables the control of spins and magnetization without dissipation Levitov et al. (1985); Edelstein (1995, 2005)—a critical functionality in realizing low-power spintronics Linder and Robinson (2015); Eschrig (2015, 2011). Microscopically, this process occurs through the superconducting magnetoelectric effect (sometimes known as the Edelstein effect) Edelstein (1995, 2005); Fujimoto (2005); Tkachov (2017); He and Law (2020, 2021); Ikeda and Yanase (2020); Chirolli et al. (2022); He et al. (2019); Yuan (2023), where a supercurrent flowing through a noncentrosymmetric metal can induce a net magnetization of the carriers. In Rashba superconductors Bauer and Sigrist (2012); Smidman et al. (2017); Houzet and Meyer (2015), such magnetoelectric effects can be naturally understood through the spin-orbit interaction (SOI): spin readily couples to the motion of electrons (and their associated Cooper pairs) producing a net spin polarization in the presence of supercurrent flow.
Major attention has focused on the linear superconducting magnetoelectric effect, where a spin response develops by driving supercurrent in time-reversal ($\mathcal{T}$) invariant noncentrosymmetric superconductors. For example, superconductors with a polar axis $\bm{c}$ and Rashba SOI Edelstein (1995, 2005) possess a supercurrent-induced magnetization $\bm{M}\propto\bm{c}\times\bm{J}_{s}$ that lies in-plane. Here $\bm{J}_{s}$ is the supercurrent density. However, the nonlinear magnetoelectric effect in superconductors is much less studied. Indeed, even in the normal state, nonlinear spin generation is a topic of intense recent interests Hamamoto et al. (2017); Xiao et al. (2022, 2023); Guo et al. (2024); Feng et al. (2024).
Figure 1: (a) Schematic illustration of the ALM/SC heterostructure exhibiting the nonlinear superconducting magnetoelectric (NSM) effect. In the presence of a supercurrent $J_{s}$ flowing along $x$ direction, the net magnetization along $z$ direction $M_{z}$ can be created. (b) In noncentrosymmetric superconductors with Rashba SOI, an in-plane magnetization, denoted as $M_{x,y}$, that is linear with $J_{s}$ can be generated. In ALM/SC heterostructures, an out-of-plane $M_{z}$ is generated that goes as $J_{s}^{2}$. When $J_{s}=0$, $M_{z}=0$ in the ALM/SC heterostrucure vanishes due to $C_{4}\mathcal{T}$ symmetry. (c) The normal-state band structure of the altermagnet from Eq. (8) at $J_{\mathrm{ex}}=0.4t$. The spin-split bands for spin $\uparrow$ and spin $\downarrow$ are indicated in red and blue respectively.
In this work, we unveil the nonlinear superconducting magnetoelectric effect (NSM) by examining second-order spin generation in response to the supercurrent. Specifically, we find a (second-order nonlinear) supercurrent-induced magnetization $\delta\bm{M}$ in two dimensions as
$\delta M_{c}^{(2)}=\chi_{ab}^{c}q_{a}q_{b},$ (1)
where $a,b=x,y$ is the direction of the supercurrent, $c=x,y,z$, and $\bm{q}$ is the Cooper pair momentum in the presence of supercurrent flow. Note $\chi_{ab}^{c}$ vanishes in the presence of $\mathcal{T}$ symmetry. As a result, to realize the second-order nonlinear response, $\mathcal{T}$ symmetry breaking is necessary. Interestingly, Eq. (1) persists even in centrosymmetric systems in sharp contrast to the case of the linear magnetoelectric effect in noncentrosymmetric superconductors He and Law (2020). As we argue, NSM can be naturally realized in magnet/superconductor heterostructures by utilizing the proximity effect between a magnetic material and an $s$-wave superconductor Fu and Kane (2008); Qi et al. (2010); Wang et al. (2015); Tkachov (2017).
For concreteness, we propose an altermagnet/superconductor (ALM/SC) heterostructure as depicted in Fig. 1(a) as a natural candidate for NSM. Altermagnets are collinear antiferromagnets with unconventional magnetic order Ma et al. (2021); Šmejkal et al. (2022a); Lee et al. (2024); Mazin et al. (2022); Mazin (2023); Feng et al. (2022); Šmejkal et al. (2022b, 2023); Ghorashi et al. (2023); Fang et al. (2023); Krempaskỳ et al. (2024); recently, several potential altermagnetic materials have been identified by ab initio simulations Ma et al. (2021); Šmejkal et al. (2022a); Lee et al. (2024). A characteristic property of $d$-wave altermagnets is $C_{4}\mathcal{T}$ symmetry, which breaks $\mathcal{T}$ and four-fold rotational ($C_{4}$) symmetry individually but preserve the combination of them. The interplay between the altermagnets and superconductivity has been theoretically investigated Zhu et al. (2023); Beenakker and Vakhtel (2023); Zhang et al. (2024); Ouassou et al. (2023); Papaj (2023); Li and Liu (2023); Giil et al. (2024).
For ALM/SC heterostructures, we find that a nonvanishing out-of-plane spin magnetization can be generated in the second-order response to the driving supercurrent. This is in sharp contrast to the case of polar or gyrotropic superconductors Edelstein (1995); He et al. (2019); He and Law (2020). This work is organized as follows: First we present a general theory of supercurrent-induced spin generation including both the linear (first order) and nonlinear (second order) terms. Then we focus on the case in a specific ALM/SC realization. Remarkably, due to $C_{4}\mathcal{T}$ symmetry in altermagnets, we find that the second-order response $\chi_{ab}^{z}$ provides the leading order contribution to the out-of-plane spin magnetization; both equilibrium out-of-plane magnetization and linear-order out-of-plane magnetoelectric effects vanish. Additionally, due to the presence of a weak SOI in real altermagnets, the linear magnetoelectric effect also shows up in the in-plane direction. Furthermore, in the weak oscillating regime, driving an AC supercurrent can induce the second harmonic and rectified magnetization. Finally, we discuss possible material candidates that include thin films of the altermagnet RuO2 and KRu4O8 on an $s$-wave superconductor. We propose that ALM/SC devices allow for the non-dissipative electric control of magnetization, which is important for high density magnetic memories and offers tantalizing possibilities for spin transport in which Joule heating and dissipation are minimized.
Superconducting magnetoelectric effect—We first examine the superconducting magnetoelectric effect within an effective Bogoliubov-de Gennes (BdG) framework. We note that our treatment is general and is agnostic to the precise microscopics of the superconducting gap: the superconducting state can be intrinsic or proximity induced. In what follows, we describe either case phenomenologically via the pairing potential $\Delta_{\bm{k}}$. The finite-$\bm{q}$ BdG Hamiltonian in the Nambu basis $(\hat{c}_{\bm{k}+\bm{q}/2,\uparrow},\hat{c}_{\bm{k}+\bm{q}/2,\downarrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\uparrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\downarrow})$ reads
$\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}=\left(\begin{matrix}{}H_{\bm{k}+\bm{q}/2}&\hat{\Delta}_{\bm{k}}\\ \hat{\Delta}^{\dagger}_{\bm{k}}&-H^{*}_{-\bm{k}+\bm{q}/2}\end{matrix}\right),$ (2)
where $H_{\bm{k}}$ is the Bloch Hamiltonian of the normal state, and $\bm{q}$ is the momentum of the Cooper pair. As we will see below, altermagnetic $H_{\bm{k}}$ enables the NSM effect. For simplicity we have used a $\bm{q}$-independent gap function $\hat{\Delta}_{\bm{k}}$, which is valid for weak values of the applied current Chirolli et al. (2022).
Table 1: Symmetry restrictions for three components of $\bm{M}$ in a two-dimensional system. Note that $C_{4}\mathcal{T}$ corresponds to the case for altermagnets that we focus on in our work. 1(2) indicates first (second) order response. $\checkmark$ indicates allowed and $\times$ indicates forbidden.
$\mathcal{P}$ $\mathcal{T}$ $\mathcal{PT}$ $C_{4}\mathcal{T}$
$M_{x}$(2) ✓(1) $\times$ ✓(1)
$M_{y}$(2) ✓(1) $\times$ ✓(1)
$M_{z}$(2) ✓(1) $\times$ ✓(2)
The total free energy $\mathcal{F}$ of the ALM/SC heterostructure comprises two parts: the superconductor component $\mathcal{F}_{\mathrm{sc}}$ and the material (i.e. altermagnet) component $\mathcal{F}_{\mathrm{al}}$ so that $\mathcal{F}=\mathcal{F}_{\mathrm{sc}}+\mathcal{F}_{\mathrm{al}}$. The supercurrent can be tracked as $\bm{J}_{s}=\nu_{s}\bm{q}$ with the condensate superfluid stiffness $\nu_{s}=\partial^{2}\mathcal{F}_{\mathrm{sc}}(\bm{q})/\partial^{2}\bm{q}$. As a result, observables that scale with $q$ below are induced by the supercurrent at linear order while $q^{2}$ indicates a second-order nonlinear response.
The magnetic response of the heterostructure can be obtained in the standard fashion by analyzing the set of parametric partition function $\mathcal{Z}(\bm{h})$
$\mathcal{Z}(\bm{h})=e^{-\beta\{\mathcal{F}_{\mathrm{sc}}-\frac{1}{2\beta}\sum_{\bm{k},\bm{q},n}\mathrm{Tr}[\mathrm{log}G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})]\}},$ (3)
where $\bm{h}$ is a parameter describing an auxiliary Zeeman field with a corresponding Zeeman energy $\Sigma(\bm{h})=\mathrm{diag}(g_{s}\mu_{B}\bm{h}\cdot\hat{s},-g_{s}\mu_{B}\bm{h}\cdot\hat{s}^{*})/2$. Here $g_{s}=2$ is the Landé $g$ factor, $\mu_{B}$ being the Bohr magneton, and the Gor’kov’s Green’s functions are $G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})=\mathcal{G}^{-1}(\bm{k},\bm{q},i\omega_{n})+\Sigma(\bm{h})$ and $\mathcal{G}(\bm{k},\bm{q},i\omega_{n})=(i\omega_{n}-\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}})^{-1}$ with the Matsubara frequency $\omega_{n}=(2n+1)\pi/\beta$. $\beta=1/k_{B}T$ with $T$ being the temperature.
The spin magnetization can be evaluated as $\bm{M}=-\beta^{-1}\partial\mathrm{log}\mathcal{Z}(\bm{h})/\partial{\bm{h}}|_{\bm{h}=0}$. Using the BdG Hamiltonian [Eq. (2)] in the current-carrying state, we obtain the spin magnetization as
$M_{a}=-\frac{g_{s}\mu_{B}}{4\beta}\sum_{n\bm{k}}\mathrm{Tr}[\mathcal{G}(\bm{k},\bm{q},i\omega_{n})\eta_{a}].$ (4)
Here $\eta_{a}=\mathrm{diag}(s_{a},-s_{a}^{*})$ are the generalized spin Pauli matrices in the Nambu space. To extract the linear and nonlinear responses systematically, we expand the BdG Hamiltonian as
$\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}=\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}=0}+\frac{1}{2}q_{a}\hat{v}_{a}+\frac{1}{8}q_{a}q_{b}\hat{w}_{ab},$ (5)
where $\hat{v}_{a}=\mathrm{diag}[\mathcal{V}_{a}(\bm{k}),-\mathcal{V}_{a}^{*}(-\bm{k})]$ and $\hat{w}_{ab}=\mathrm{diag}[\mathcal{W}_{ab}(\bm{k}),-\mathcal{W}_{ab}^{*}(-\bm{k})]$ with $\mathcal{V}_{a}(\bm{k})=\partial H_{\bm{k}}/\partial k_{a}$ and $\mathcal{W}_{ab}(\bm{k})=\partial^{2}H_{\bm{k}}/\partial k_{a}\partial k_{b}$. At linear order in $\bm{q}$, we find $\delta M_{a}^{(1)}=\alpha_{ab}q_{b}$, where the first-order spin susceptibility $\alpha_{ab}$ reads
$\alpha_{ab}=-\frac{g_{s}\mu_{B}}{8\beta}\sum_{n\bm{k}}\mathrm{Tr}[\eta_{a}\mathcal{G}_{0}\hat{v}_{b}\mathcal{G}_{0}],$ (6)
where $\mathcal{G}_{0}\equiv\mathcal{G}(\bm{k},0,i\omega_{n})=(i\omega_{n}-\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}=0})^{-1}$ for brevity. This formula is consistent with the results in Ref. Chirolli et al. (2022). Importantly, we find the nonlinear spin magnetization along $c$ direction in Eq. (1) with the second-order spin susceptibility $\chi$ as
$\chi_{ab}^{c}=-\frac{g_{s}\mu_{B}}{32\beta}\!\sum_{n\bm{k}}\!\mathrm{Tr}[\eta_{c}(\mathcal{G}_{0}\hat{w}_{ab}\mathcal{G}_{0}+2\mathcal{G}_{0}\hat{v}_{a}\mathcal{G}_{0}\hat{v}_{b}\mathcal{G}_{0})].$ (7)
We note that that $\chi_{ab}^{c}$ in Eq. (7) is general for arbitrary model Bloch Hamiltonians as well as pairing potentials. In particular, Eq. (7) applies for both conventional pairing as well as unconventional pairing $\hat{\Delta}_{\bm{k}}$. However, in the next section we will take a simple $s$-wave pairing $\hat{\Delta}_{\bm{k}}=\Delta_{0}is_{y}$ with a $\bm{k}$-independent pairing gap as an illustrative example to emphasize the effect of the magnetic order. Detailed derivations can be found in Supplementary Material Not .
It is instructive to examine the symmetry requirements for both the linear $\alpha_{ab}$ and second-order NSM $\chi_{ab}^{c}$ summarized in Table 1. For instance, $\mathcal{T}$ breaking is necessary for non-zero $\chi_{ab}^{c}$ while $\mathcal{P}$ breaking is necessary for $\alpha_{ab}$; both are destroyed in $\mathcal{PT}$-invariant systems. A particular interesting case is that of $C_{4}\mathcal{T}$ symmetry found in altermagnets: we find that $\chi_{ab}^{x}$ and $\chi_{ab}^{y}$ terms vanish; only $\chi_{ab}^{z}$ is non-zero. Interestingly, $C_{4}\mathcal{T}$ further constrains the NSM so that $\chi_{xx}^{z}=-\chi_{yy}^{z}$. In the presence of a weak SOI naturally found in realistic altermagnets, in-plane spin magnetization can be induced as expected of a linear magnetoelectric effect in Eq. (6). Importantly, however, the Rashba SOI induced linear magnetoelectric effect is purely in-plane ($\alpha_{zx}=\alpha_{zy}=0$). This means that for out-of-plane magnetatization responses, the second order NSM $\chi_{ab}^{z}$ is the leading order term in the altermagnetic systems. As we will see below, it dominates the out-of-plane magnetization responses.
Figure 2: (a) The fully gapped BdG energy spectrum of ALM/SC from Eq. (2) with Bloch Hamilotnian in Eq. (8) at $k_{y}=0$ using a phenomelogical gap $\hat{\Delta}_{\bm{k}}=\Delta_{0}is_{y}$. (b) The spin magnetization $M_{z}$ as a function of $q_{x}$ from Eq. (4) (solid orange). As a comparison, the NSM susceptibility multiplied by $q_{x}^{2}$ is plotted as Eq. (7) (dashed green). (c) The nonlinear spin susceptibility $\chi^{z}_{xx}$ as a function of $d$-wave magnetic order $J_{\mathrm{ex}}$. The inset shows the region of $0<J_{\mathrm{ex}}<0.04t$. (d) For nonzero $\lambda$, the linear and nonlinear spin susceptibilities $\alpha_{xy}$ and $\chi^{z}_{xx}$ as a function of $\lambda$. $\alpha_{xy}$ increases linearly with $\lambda$ for small $\lambda$. Parameters: $\lambda=0$ for (a), (b) and (c). $J_{\mathrm{ex}}=0.2t$ for (a), (b) and (d). $(\Delta_{0},\mu)=(0.1,0.3)t$ for all four panels. The temperature is set to be $T=0.3T_{c}$.
NSM in ALM/SC heterostructure—In the set-up illustrated in Fig. 1(a), an altermagnet thin film is in contact with a conventional $s$-wave superconductor. To characterize the physics of quasi-2D planar $d_{x^{2}-y^{2}}$-wave altermagnet, we adopt a generic two-band Hamiltonian Šmejkal et al. (2022a), yielding
$H_{\bm{k}}=[ta^{2}(k_{x}^{2}+k_{y}^{2})-\mu]s_{0}+J_{\mathrm{ex}}a^{2}(k_{x}^{2}-k_{y}^{2})s_{z}.$ (8)
Here $t$ parameterizes the usual kinetic energy, $J_{\mathrm{ex}}$ denotes the $d_{x^{2}-y^{2}}$ exchange magnetic order and $\mu$ is the Fermi energy. $a$ denotes the lattice constant. We note that an additional Rashba SOI term $H_{R}=\lambda a(k_{x}s_{y}-k_{y}s_{x})$ can be readily added to $H_{\bm{k}}$ with $\lambda$ term representing the strength of SOI, that arises in altermagnets Šmejkal et al. (2022a); Fang et al. (2023). Here, $s$ matrices operate on the spin degree of freedom. This Hamiltonian breaks time reversal symmetry $\mathcal{T}=-is_{y}\mathcal{K}$ ($\mathcal{K}$ is the complex conjugate) and four-fold rotation symmetry $C_{4}=e^{i\pi s_{z}/4}$, but preserves the combination of them as $C_{4}\mathcal{T}$. The SOI term breaks inversion ($\mathcal{P}$) symmetry without breaking $C_{4}\mathcal{T}$. Throughout our work, we use the energy unit $t$. As an illustration of the spin-split band structure in altermagnet, we plot Eq. (8) in Fig. 1(c); here we have used $J_{\mathrm{ex}}=0.4t$ and $\lambda=0$.
Because of the proximity effect between the altermagnet and superconductor, Cooper pairs can tunnel into the altermagnet. For a thin film of altermagnet [see Fig. 1(a)], with thickness $d$ much less than the coherence length of the superconductor ($d\ll\xi$), the pairing gap is approximately uniform along the $z$ direction. We have focused on an conventional $s$-wave superconductor with the pairing $\Delta_{0}$. Using the parameters as $(J_{\mathrm{ex}},\lambda,\mu,\Delta_{0})=(0.2,0,0.3,0.1)t$, we diagonalize the BdG Hamiltonian and plot the fully gapped BdG energy spectrum in Fig. 2(a). Although the pairing potential is purely $s$-wave, the induced superconducting correlations can be obtained as $F=[\psi_{0}+\bm{d}(\bm{k})\cdot\bm{s}]is_{y}$, where $\psi_{0}$ and $\bm{d}$ vector parameterize the spin-singlet and spin-triplet pairing correlations, respectively Gor’kov and Rashba (2001); Frigeri et al. (2004); Zhou et al. (2016). We have $d=(0,0,d_{z})$ with $d_{z}\propto J_{\mathrm{ex}}(k_{x}^{2}-k_{y}^{2})$, indicating that the $d$-wave magnetic order gives rise to the spin-triplet pairing correlation in altermagnets due to the proximity effect.
To illustrate NSM effect in the ALM/SC heterostructure, we numerically study the supercurrent induced magnetization in Fig. 2 for the Bloch Hamiltonian in Eq. (8). By applying the current along $x$ direction as $\bm{q}=(q_{x},0)$, we calculate the magnetization directly from Eq. (4) at $J_{\mathrm{ex}}=0.2t$ as shown in Fig. 2(b). One finds that $M_{z}\sim q_{x}^{2}$ as expected from Fig. 1(b). It is also worth noting that when $q_{x}=0$, equilibrium magnetization vanishes $M_{z}=0$, which is consistent with $C_{4}\mathcal{T}$ symmetry in altermagnets. We note that numerically computing the NSM susceptibility $\chi_{xx}^{z}$ from Eq. (7) and multiplying by $q_{x}^{2}$ as plotted in the dashed green line matches the solid line in orange [from Eq. (4)].
In Fig. 2(c), we show the second-order spin susceptibility $\chi_{xx}^{z}$ as a function of $J_{\mathrm{ex}}$ from Eq. (7). Generally, $\chi_{xx}^{z}$ starts to grow linearly with $J_{\mathrm{ex}}$ as shown in the zoom-in inset of Fig. 2(c). Furthermore, we also calculate the first-order spin susceptibility $\alpha_{xy}$ (purple) which is shown in Fig. 2(d). As expected, $\alpha_{xy}$ directly depends on $\lambda$; for small $\lambda$, $\alpha_{xy}$ increases linearly with $\lambda$ and vanishes when $\lambda=0$. In contrast, for NSM $\chi_{xx}^{z}$ is finite for $\lambda=0$ and exhibits a weak dependence on $\lambda$; $\chi_{xx}^{z}$ becomes smaller when $\lambda$ is nonzero since Rashba SOI tends to pin the spin in the in-plane direction and weakens the out-of-plane spin polarization. In our calculations, we adopt the BCS temperature dependence (for the $s$-wave superconductor substrate) of $\Delta_{0}$ with $\Delta(T)=\Delta_{0}\tanh(1.74\sqrt{T_{c}/T-1})$ and $\Delta_{0}=1.76k_{B}T_{c}$.
Figure 3: The second-order spin susceptibility $\chi_{xx}^{z}$ as a function of the temperature $T$ at $(J_{\mathrm{ex}},\Delta_{0},\mu)=(0.02,0.1,0.3)t$. The purple diamonds denotes the numerical results from Eq. (7) and the orange line denotes the analytical result from Eq. (9). At low temperatures, $\chi^{z}_{xx}\sim\beta^{3}\mathrm{sech}^{2}\beta$.
[…]
$\begin{split}f(x)=&\frac{\beta^{3}J_{\mathrm{ex}}}{64\pi}x\mathrm{sech}^{2}\gamma\Big{[}\frac{4}{\beta^{2}}+t^{2}x^{4}(2-3\mathrm{sech}^{2}\gamma)\\ &+\frac{4tx^{2}(\mu-tx^{2})\tanh\gamma}{\gamma}\Big{]},\end{split}$ (9)
where $\gamma=\beta\sqrt{\Delta(T)^{2}+(tx^{2}-\mu)^{2}}/2$. By employing Eq. (9), we plot the $\chi_{xx}^{z}$ as a function of temperature in Fig. 3 (orange solid curve) with $J_{\mathrm{ex}}=0.02t$. By way of comparison, we also show a fully numerical plot of Eq. (7) (purple diamonds); both agree with each other. Interestingly, $\chi_{xx}^{z}$ vanishes at zero temperature. At low temperatures ($\beta\Delta_{0}\gg 1$), $\chi_{xx}^{z}$ starts to grows with $T$ scaling as $\chi_{xx}\sim\beta^{3}\mathrm{sech}^{2}\beta$. However, at $T/T_{c}\sim 0.6$, it reaches a peak after which it rapidly diminishes with temperature. This non-monotonic behavior contrasts sharply with that of the linear magnetoelectric effect in Rashba-type superconductors He et al. (2019); Not that instead saturates at low temperature.
In superconductors, supercurrent flow can arise due to either an external applied magnetic field or an explicit transport current Anthore et al. (2003). In the former case, there is a screening supercurrent on the surface of the superconductor. In the latter case, the global excitation of the superconductor gives a uniform Cooper pair momentum $\bm{q}$. Specifically, for slowly varying AC supercurrent as $\bm{J}_{s}(t)=\bm{J}_{\omega}\cos\omega t$ with $\omega\ll\Delta_{0}$, we expect an oscillatory Cooper pair momentum $\bm{q}(t)=\bm{q}_{\omega}\cos\omega t$. This will allow to obtain a second harmonic magnetization in the weak oscillating regime with $\chi_{ab}^{c}(2\omega)=\chi_{ab}^{c}/4$.
Candidate materials.—We anticipate that NSM we discuss here can be realized in ALM/SC formed out of readily available $d$-wave altermagnetic materials Šmejkal et al. (2022a). As an example, consider the altermagnet RuO2 with in-plane lattice constant $a=4.5\AA$, and $t=2.5$ eV, $J_{\mathrm{ex}}=0.5$ eV Šmejkal et al. (2022b); Zhang et al. (2024). Using a pairing gap of $\Delta_{0}=1$ meV, $\mu=10$ meV, $\lambda=10$ meV and $T=0.3T_{c}$ we estimate $\chi_{xx}^{z}\approx 0.07\mu_{B}$. For a Cooper pair momentum $q\approx 10^{-3}\AA^{-1}$ Zhu et al. (2021), the induced out-of-plane spin density in the RuO2 thin film/superconductor heterostructure is $\sim 0.7\times 10^{-5}\mu_{B}/\mathrm{nm}^{2}$. Note that this value is comparable to the spin generation in MnBi2Te4 and other noncentrosymmetric ferromagnetic systems Xiao et al. (2022); Chernyshov et al. (2009); Fang et al. (2011). Another example is KRu4O8 with the parameter values: $t=0.05$ eV, $J_{\mathrm{ex}}=0.018$ eV, and $a=9.9\AA$. We find $\chi_{xx}^{z}\approx 0.01\mu_{B}$ and out-of-plane spin density similar to that discussed above. These magnitudes are estimated qualitatively using a low-energy effective description of altermagnets in the small $J$ and weak doping (small $\mu$) regime as well as a gapped BdG spectrum. In real altermagnets with large $J$ and high doping, the BdG spectrum becomes gapless. We anticipate this may enhance NSM with additional contribution from the quasiparticles excited at finite temperature.
In this work, we have proposed a novel nonlinear superconducting magnetoelectric effect. Naturally occurring in ALM/SC heterostructures, this second-order spin generation in response to the supercurrent becomes the leading order contribution to the magnetization along the out-of-plane direction. Supercurrent induced magnetization can be readily detected by using a superconducting quantum interference device (SQUID) perpendicular to the altermagnet surface, which can be employed to probe the magnetic flux change Chirolli et al. (2022). Interestingly, nonlinear spin generation that is second order in an applied electric field can be realized in normal-state (non-superconducting) altermagnets with $C_{4}\mathcal{T}$ symmetry Xiao et al. (2022). Our work shows that nonlinear spin generation survives in the superconducting state and is driven by a dissipationless supercurrent.
Acknowledgements.—We thank Bo Yang and Wen-Yu He for helpful discussions. This work was supported by the Ministry of Education Singapore under its Academic Research Fund Tier 2 Grant No. MOE-T2EP50222-0011.
Note added.—Recently, we became aware of an independent pre-print posted on the arXiv in Ref. Zyuzin (2024) which also found a second-order supercurrent-induced spin polarization in superconductors with $d$-wave magnetization.
Supplementary Material for
“Nonlinear Superconducting Magnetoelectric Effect”
Jin-Xin Hu,1 Oles Matsyshyn,1 Justin C. W. Song1
1Division of Physics and Applied Physics, School of Physical and Mathematical Sciences, Nanyang Technological University, Singapore 637371
I I. Theory of nonlinear superconducting Magnetoelectric effect
I.1 A. Green’s function method
In this section we provide more details for deriving the general formula of superconducting magnetoelectric effect, including both the linear and nonlinear terms. The superconducting state can be intrinsic or proximity induced with a pairing function $\hat{\Delta}_{\bm{k}}$. $\hat{\Delta}_{\bm{k}}$ can be conventional $s$-wave with a uniform pairing gap $\hat{\Delta}_{\bm{k}}=\Delta_{0}is_{y}$ or unconventional pairing states. In the current-carrying state, the finite-$\bm{q}$ Bogoliubov-de Gennes (BdG) Hamiltonian in the Nambu basis $(\hat{c}_{\bm{k}+\bm{q}/2,\uparrow},\hat{c}_{\bm{k}+\bm{q}/2,\downarrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\uparrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\downarrow})$ reads
$\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}=\left(\begin{matrix}{}H_{\bm{k}+\bm{q}/2}&\hat{\Delta}_{\bm{k}}\\ \hat{\Delta}^{\dagger}_{\bm{k}}&-H^{*}_{-\bm{k}+\bm{q}/2}\end{matrix}\right)$ (S1)
where $H_{\bm{k}}$ is the Bloch Hamiltonian of the normal state, $\bm{q}$ is the momentum of the Cooper pair, and $s_{y}$ is a Pauli matrix. For simplicity we have used a $\bm{q}$-independent gap function $\hat{\Delta}_{\bm{k}}$, which is valid for weak values of the applied current.
In the main text, we focus on the altermagnet/superconductor heterostructure. The total free energy $\mathcal{F}$ of the system contains the superconductor part $\mathcal{F}_{\mathrm{sc}}$ and the altermagnet part $\mathcal{F}_{\mathrm{al}}$ with $\mathcal{F}=\mathcal{F}_{\mathrm{sc}}+\mathcal{F}_{\mathrm{al}}$. The supercurrent $J_{s}=\nu_{s}\bm{q}$ with the superfluid density $\nu_{s}=\partial^{2}\mathcal{F}_{\mathrm{sc}}(\bm{q})/\partial^{2}\bm{q}$. The partition function $\mathcal{Z}(\bm{h})$ is given by
$\mathcal{Z}(\bm{h})=e^{-\beta[\mathcal{F}_{\mathrm{sc}}-\frac{1}{2\beta}\sum_{\bm{k},\bm{q},n}\mathrm{Tr}[\mathrm{log}G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})]]}.$ (S2)
By introducing an auxiliary Zeeman field $\bm{h}$, we can write down the modified BdG Hamiltonian:
$\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}(\bm{h})=\left(\begin{array}[]{cc}H_{\bm{k}+\bm{q}/2}-\frac{1}{2}g_{s}\mu_{B}\bm{h}\cdot\bm{s}&\hat{\Delta}_{\bm{k}}\\ \hat{\Delta}^{\dagger}_{\bm{k}}&-H^{*}_{-\bm{k}+\bm{q}/2}+\frac{1}{2}g_{s}\mu_{B}\bm{h}\cdot\bm{s}^{*}\end{array}\right),$ (S3)
This auxiliary Zeeman field is to be distinguished from the genuine magnetization of the system and is set to zero at the end of the calculation. The Gor’kov’s Green’s function is $G(\bm{k},\bm{q},\bm{h},i\omega_{n})=[i\omega_{n}-\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}(\bm{h})]^{-1}$. The magnetization $M$ can be obtained from the partition function as
$\begin{split}M_{a}&=-\frac{1}{\beta}\frac{\partial}{\partial{h_{a}}}\mathrm{log}Z(\bm{h})\\ &=-\frac{1}{2\beta}\sum_{\bm{k},\bm{q},n}\frac{\partial}{\partial{h_{a}}}\mathrm{Tr}[\mathrm{log}G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})]\end{split}$ (S4)
We decompose the Green’s function as $G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})=\mathcal{G}^{-1}(\bm{k},\bm{q},i\omega_{n})+\Sigma(\bm{h})$ with
$\Sigma(\bm{h})=\left(\begin{array}[]{cc}\frac{1}{2}g_{s}\mu_{B}\bm{h}\cdot\bm{s}&0\\ 0&-\frac{1}{2}g_{s}\mu_{B}\bm{h}\cdot\bm{s}\end{array}\right),$ (S5)
We can then expand the logarithm in the first order of $\bm{h}$ as $\mathrm{log}G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})=\mathrm{log}\mathcal{G}^{-1}(\bm{k},\bm{q},i\omega_{n})+\mathcal{G}(\bm{k},\bm{q},i\omega_{n})\Sigma(\bm{h})$. Thus, we can then obtain the magnetization $M$ as
$M_{a}=-\frac{g_{s}\mu_{B}}{4\beta}\sum_{n\bm{k}}\mathrm{Tr}[\mathcal{G}(\bm{k},\bm{q},i\omega_{n})\eta_{a}].$ (S6)
Here $\eta_{a}=\mathrm{diag}(s_{a},-s_{a}^{*})$ is the redefined spin Pauli matrices in the Nambu space. The Gor’kov Green’s function is $\mathcal{G}(\bm{k},\bm{q},i\omega_{n})=(i\omega_{n}-\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}})^{-1}$ with the Matsubara frequency $\omega_{n}=(2n+1)\pi k_{B}T$. $T$ is the temperature. $g_{s}=2$ is the Landé $g$ factor and $\mu_{B}$ is the Bohr magneton. We can then expand the BdG Hamiltonian as
$\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}=\mathcal{H}_{\mathrm{BdG}}^{\bm{k}}+q_{a}\hat{v}_{a}/2+q_{a}q_{b}\hat{w}_{ab}/8+\mathcal{O}(\bm{q}^{3})$ (S7)
The velocity operators are
$\hat{v}_{a}=\left(\begin{array}[]{cc}\hat{\mathcal{V}}_{a}(\bm{k})&0\\ 0&-\hat{\mathcal{V}}^{*}_{a}(-\bm{k})\end{array}\right),\hat{w}_{ab}=\left(\begin{array}[]{cc}\mathcal{W}_{ab}(\bm{k})&0\\ 0&-\mathcal{W}_{ab}^{*}(-\bm{k})\end{array}\right).$ (S8)
Here we introduce $\mathcal{V}_{a}(\bm{k})=\partial H_{\bm{k}}/\partial k_{a}$ and $\mathcal{W}_{ab}(\bm{k})=\partial^{2}H_{\bm{k}}/\partial k_{a}\partial k_{b}$. We proceed in evaluating terms of Eq. (S7), we note that if $A$ and $B$ are matrices and $B\ll A$ ($B\sim\mathcal{O}(\bm{q})$), we can expand the matrix expression in small $B$ as
$(A+B)^{-1}=A^{-1}-A^{-1}BA^{-1}+A^{-1}BA^{-1}BA^{-1}+\mathcal{O}(B^{3}).$ (S9)
In the following we will apply Eq. (S9) to the full Green function in order to make an effective expansion around small $\bm{q}$. We obtain
$\mathcal{G}(\bm{k},\bm{q},i\omega_{n})=(i\omega_{n}-\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}})^{-1}=\mathcal{G}_{0}+\frac{1}{2}q_{a}\mathcal{G}_{0}\hat{v}_{a}\mathcal{G}_{0}+q_{a}q_{b}(\frac{1}{8}\mathcal{G}_{0}\hat{w}_{ab}\mathcal{G}_{0}+\frac{1}{4}\mathcal{G}_{0}\hat{v}_{a}\mathcal{G}_{0}\hat{v}_{b}\mathcal{G}_{0})$ (S10)
Here $\mathcal{G}_{0}\equiv\mathcal{G}(\bm{k},i\omega_{n})=(i\omega_{n}-\mathcal{H}_{\mathrm{BdG}}^{\bm{k}})^{-1}$. We insert Eq. (S10) to Eq. (S6) to get the magnetization susceptibility at both linear- and (nonlinear) second-order. Firstly, the first-order spin susceptibility $\alpha_{ab}$ is described by $\delta M_{a}^{(1)}=\alpha_{ab}q_{b}$, where $\alpha_{ab}$ reads
$\alpha_{ab}=-\frac{g_{s}\mu_{B}k_{B}T}{8}\sum_{n\bm{k}}\mathrm{Tr}[\eta_{a}\mathcal{G}_{0}\hat{v}_{b}\mathcal{G}_{0}].$ (S11)
More importantly, we can write the second-order nonlinear spin magnetization along $c$ direction as $\delta M_{c}^{(2)}=\chi_{ab}^{c}q_{a}q_{b}$, and the nonlinaer susceptibility $\chi$ can be obtained as
$\chi_{ab}^{c}=-\frac{g_{s}\mu_{B}k_{B}T}{32}\!\sum_{n\bm{k}}\!\mathrm{Tr}[\eta_{c}(\mathcal{G}_{0}\hat{w}_{ab}\mathcal{G}_{0}+2\mathcal{G}_{0}\hat{v}_{a}\mathcal{G}_{0}\hat{v}_{b}\mathcal{G}_{0})].$ (S12)
I.2 B. Analytical result
In the previous subsection, we use the Green’s function method to give the general formulas of the linear and nonlinear magnetization. These formulas can be adopt to evaluate the spin susceptibilities for altermagnet/superconductor heterostructures. In this subsection, we provide the analytical solution of the nonlinear spin susceptibility $\chi^{z}_{ab}$. For the altermagnets, we can write down a general two-band model
$H_{\bm{k}}=[ta^{2}(k_{x}^{2}+k_{y}^{2})-\mu]s_{0}+J_{ex}a^{2}[2k_{x}k_{y}\cos(2\theta)+(k_{x}^{2}-k_{y}^{2})\sin(2\theta)]s_{z}$ (S13)
$t$ parameterizes the usual kinetic energy, $J_{ex}$ denotes the $d$-wave exchange magnetic order parameter and $\mu$ is the chemical potential. $\theta$ is the angle between the altermagnetic orientation and the crystalline axes. The velocity operators are
$\hat{v}_{x}=a^{2}\left(\begin{array}[]{cc}2tk_{x}s_{0}+2J_{ex}k\sin(\phi+2\theta)s_{z}&0\\ 0&2tk_{x}s_{0}+2J_{ex}k\sin(\phi+2\theta)s_{z}\end{array}\right),$ (S14)
$\hat{v}_{y}=a^{2}\left(\begin{array}[]{cc}2tk_{y}s_{0}+2J_{ex}k\cos(\phi+2\theta)s_{z}&0\\ 0&2tk_{y}s_{0}+2J_{ex}k\cos(\phi+2\theta)s_{z}\end{array}\right),$ (S15)
and
$\hat{w}_{xx}=a^{2}\left(\begin{array}[]{cc}2ts_{0}+2J_{ex}\sin(2\theta)s_{z}&0\\ 0&-2ts_{0}-2J_{ex}\sin(2\theta)s_{z}\end{array}\right),\hat{w}_{xy}=a^{2}\left(\begin{array}[]{cc}2J_{ex}\cos(2\theta)s_{z}&0\\ 0&-2J_{ex}\cos(2\theta)s_{z}\end{array}\right),$ (S16)
[…]
$f(x)=\frac{\beta^{3}J_{\mathrm{ex}}}{64\pi}x\mathrm{sech}^{2}\gamma\Big{[}\frac{4}{\beta^{2}}+t^{2}x^{4}(2-3\mathrm{sech}^{2}\gamma)+\frac{4tx^{2}(\mu-tx^{2})\tanh\gamma}{\gamma}\Big{]},$ (S17)
where $\gamma=\beta\sqrt{\Delta^{2}+(tx^{2}-\mu)^{2}}/2$.
II II. Pairing correlations
Although we only consider the simplest conventional spin singlet pairing, i.e., $\Delta_{0}is_{y}$, here we show that the $d$-wave magnetic order will induce the spin-triplet correlations. To show this, the BdG Hamiltonian is
$\mathcal{H}_{\mathrm{BdG}}^{\bm{k}}=\begin{pmatrix}H_{\bm{k}}&\Delta_{0}is_{y}\\ (\Delta_{0}is_{y})^{\dagger}&-H^{*}_{-\bm{k}}\end{pmatrix}$ (S18)
Here $H_{\bm{k}}$ is from Eq.(S13). Let us identify the superconducting properties in terms of Green’s function:
$G_{\lambda\mu}(\bm{k},\tau)$ $=T_{\tau}\{c_{\bm{k},\lambda}(\tau)c^{\dagger}_{\bm{k},\mu}(0)\},$ (S19)
$F_{\lambda\mu}(\bm{k},\tau)$ $=T_{\tau}\{c_{\bm{k},\lambda}(\tau)c_{-\bm{k},\mu}(0)\}.$ (S20)
[…]
$\mathcal{G}(\bm{k},i\omega_{n})=(i\omega_{n}-H_{\mathrm{BdG}}^{\bm{k}})^{-1}=\begin{pmatrix}G_{e}(\bm{k},i\omega_{n})&F(\bm{k},i\omega_{n})\\ F^{\dagger}(\bm{k},i\omega_{n})&G_{h}(\bm{k},i\omega_{n})\end{pmatrix}.$ (S21)
Substitute the BdG Hamiltonian into Eq. (S21) and after some massage, we can parameterize the pairing correlation as
$F(\bm{k},i\omega_{n})=\Delta_{0}[C_{1}(\bm{k},i\omega_{n})+C_{2}(\bm{k},i\omega_{n})\bm{d}(\bm{k})\cdot\bm{s}]is_{y}$ (S22)
with the coefficients
$C_{1}(\bm{k},i\omega_{n})=-\frac{1}{2}[\frac{1}{\Delta_{0}^{2}+\xi_{k}^{2}-(J_{ex}f_{k}^{2}+i\omega_{n})^{2}}+\frac{1}{\Delta_{0}^{2}+\xi_{k}^{2}-(J_{ex}f_{k}^{2}-i\omega_{n})^{2}}],$ (S23)
$C_{2}(\bm{k},i\omega_{n})=\frac{2i\omega_{n}}{(\Delta_{0}^{2}+\xi_{k}^{2}-J_{ex}^{2}f_{k}^{2})^{2}+2(\Delta_{0}^{2}+\xi_{k}^{2}+J_{ex}^{2}f_{k}^{2})\omega_{n}^{2}+\omega_{n}^{4}},$ (S24)
where $\xi_{k}=ta^{2}(k_{x}^{2}+k_{y}^{2})-\mu$ and $f_{k}=a^{2}[2k_{x}k_{y}\cos(2\theta)+(k_{x}^{2}-k_{y}^{2})\sin(2\theta)]$. Importantly, the triplet vector is directly related to the $d$-wave magnetic order with
$\bm{d}(\bm{k})=(0,0,J_{ex}f_{k}),$ (S25)
Therefore, it can be seen that due to the presence of $d$-wave, the spin-singlet pairing and the spin-triplet pairing are mixed. Microscopically, the spin-singlet pairing correlations provide the physical origin of the nonlinear spin magnetization.
III III. Linear Edelstein effect in Rashba-type superconductors
In this section we review the linear magnetoelectric effect in superconductors with Rashba SOC. The low energy Bloch Hamiltonian is
$H_{\bm{k}}=[ta^{2}(k_{x}^{2}+k_{y}^{2})-\mu]s_{0}+\lambda a(k_{x}s_{y}-k_{y}s_{x}),$ (S26)
where $\lambda$ characterizes the strength of Rashba SOI, and $a$ is the lattice constant. We introduce the spin singlet pairing $\Delta_{\bm{k}}=\Delta_{0}is_{y}$. Similarly, the linear spin susceptibility $\alpha_{yx}$ can be obtained as
[…]
where
$\begin{split}f(x)=&\frac{\beta}{32\pi(tx^{2}-\mu)}\{x(\lambda-2tx)(tx^{2}-\mu)\mathrm{sech}^{2}\gamma_{1}+x(\lambda+2tx)(tx^{2}-\mu)\mathrm{sech}^{2}\gamma_{2}\\ &-[\Delta^{2}+(\mu-tx^{2})(\mu+\lambda x-tx^{2})]\tanh\gamma_{1}/\gamma_{1}+[\Delta^{2}+(\mu-tx^{2})(\mu-\lambda x-tx^{2})]\tanh\gamma_{2}/\gamma_{2}\}\end{split}$ (S28)
with
$\gamma_{1}=\frac{\beta\sqrt{\Delta^{2}+(\mu+\lambda x-tx^{2})^{2}}}{2},\gamma_{2}=\frac{\beta\sqrt{\Delta^{2}+(\mu-\lambda x-tx^{2})^{2}}}{2}$ (S29)
Figure S1: The linear spin magnetization in Rashba-type superconductors. (a) The band structure with Rashba SOC at $\lambda=0.3t$. (b) The first-order spin susceptibility $\alpha_{yx}$ as a function of $\lambda$ at $T=0.1T_{c}$, $\Delta_{0}=0.03t$ and $\mu=0.2t$. The purple diamonds denotes the numerical results from Eq. (S11) and the orange line denotes the analytical result from Eq. (S28). (c) The first-order spin susceptibility $\alpha_{yx}$ as a function of $\mu$ at $T=0.1T_{c}$ for various values of the pairing gap $\Delta_{0}$. (d) $\alpha_{yx}$ as a function of $T$ at $\Delta_{0}=0.1\lambda$ for various values of $\mu$.
We apply the above formula to calculate the linear (first order) magnetization in superconductors with Rashba SOI. The calculated band structure from Eq. (S26) is shown in Fig. S1(a) at $\lambda=0.3t$. By employing Eq. (S28), we plot the $\alpha_{yx}$ as a function of $\lambda$ in Fig. S1 (orange solid curve). By way of comparison, we also show a fully numerical plot of Eq. (S11) (blue diamonds); both agree with each other. In Fig. S1(c) we show the $\mu$-dependence of the $\alpha_{yx}$ with various values of $\Delta_{0}$. When $\mu<0$, $\alpha_{yx}$ decreases as $\mu$ goes down. Fig. S1(d) shows $\alpha_{yx}$ as a function of temperature for different values of $\mu$. Generally, $\alpha_{yx}$ starts to grow linearly when $T<T_{c}$, and it saturates when $T\rightarrow 0$. These behaviors are consistent with the results shown in Ref. He et al. (2019).

### v5

@@
 Abstract
-A supercurrent flow can induce a nonvanishing spin magnetization in noncentrosymmetric superconductors with spin-orbit interaction.
+Supercurrent flow can induce a nonvanishing spin magnetization in noncentrosymmetric superconductors with spin-orbit interaction.
 Often known as the non-dissipative magnetoelectric effect, these are most commonly found at linear order in supercurrent flow.
-Here, we argue that a nonlinear superconducting magnetoelectric effect (NSM) can naturally manifest in altermagnet/superconductor (ALM/SC) heterostructures: NSM manifests as a spin polarization generated as a second-order response to a driving supercurrent.
-Strikingly, we find NSM is the leading order magnetization response in ALM/SC heterostructures and survives even in the presence of centrosymmetry; $C_{4}\mathcal{T}$ symmetry in altermagnets zeroes both the equilibrium magnetization as well as out-of-plane linear magnetoelectric response.
-This renders NSM a powerful electric and non-dissipative means of controlling magnetization in ALM/SC heterostructures, a promising platform for superconducting spintronics.
+Here, we propose that a nonlinear superconducting magnetoelectric (NSM) effect can naturally manifest in magnet/superconductor heterostructures.
+In such platforms, NSM manifests as the spin polarization generated as a second-order response to a driving supercurrent.
+Strikingly, we find NSM survives centrosymmetry and is the leading order magnetic response in a variety of magnetic materials that include both collinear magnets [e.g., $d$-wave planar altermagnet thin film/superconductor (ALM/SC)] as well as noncollinear magnets (e.g., Kagome/SC systems).
+This renders NSM a powerful electric and non-dissipative means of controlling magnetization in magnet/superconductor heterostructures, a promising platform for superconducting spintronics.
 Introduction.—Unlike the normal metallic state, the intrinsically non-dissipative flow of current in superconductors enables the control of spins and magnetization without dissipation Levitov et al.
@@
 (2022); He et al.
-(2019); Yuan (2023), where a supercurrent flowing through a noncentrosymmetric metal can induce a net magnetization of the carriers.
+(2019); Yuan (2023), where a supercurrent flowing through a noncentrosymmetric metal can induce a net magnetization.
 In Rashba superconductors Bauer and Sigrist (2012); Smidman et al.
@@
 The spin-split bands for spin $\uparrow$ and spin $\downarrow$ are indicated in red and blue respectively.
-In this work, we unveil the nonlinear superconducting magnetoelectric effect (NSM) by examining second-order spin generation in response to the supercurrent.
-Specifically, we find a (second-order nonlinear) supercurrent-induced magnetization $\delta\bm{M}$ in two dimensions as
+Here we unveil the nonlinear superconducting magnetoelectric (NSM) effect by examining second-order spin generation in response to the supercurrent.
+Specifically, we find a (second-order nonlinear) supercurrent-induced magnetization $\delta\bm{M}$ in a two dimensional platform as
 $\delta M_{c}^{(2)}=\chi_{ab}^{c}q_{a}q_{b},$ (1)
@@
 Interestingly, Eq.
-(1) persists even in centrosymmetric systems in sharp contrast to the case of the linear magnetoelectric effect in noncentrosymmetric superconductors He and Law (2020).
+(1) persists in centrosymmetric systems in sharp contrast to the case of the linear magnetoelectric effect in noncentrosymmetric superconductors He and Law (2020).
 As we argue, NSM can be naturally realized in magnet/superconductor heterostructures by utilizing the proximity effect between a magnetic material and an $s$-wave superconductor Fu and Kane (2008); Qi et al.
@@
 (2015); Tkachov (2017).
-For concreteness, we propose an altermagnet/superconductor (ALM/SC) heterostructure as depicted in Fig.
-1(a) as a natural candidate for NSM.
-Altermagnets are collinear antiferromagnets with unconventional magnetic order Ma et al.
+To be specific, we propose that NSM can dominate magnetoelectricity in a variety of magnetic material/superconductor heterostructures that include both collinear (e.g., altermagnets) as well as noncollinear (e.g., kagome lattice Lee et al.
+(2024a)) magnetic order.
+A particularly striking example is $d$-wave planar altermagnet thin film/superconductor (ALM/SC) heterostructure [see Fig.
+1(a)].
+Altermagnets are collinear antiferromagnets with an unconventional magnetic order Ma et al.
 (2021); Šmejkal et al.
 (2022a); Lee et al.
-(2024); Mazin et al.
+(2024b); Mazin et al.
 (2022); Mazin (2023); Feng et al.
@@
 (2024); Fang et al.
+(2024); Sattigeri et al.
 (2023); Krempaskỳ et al.
-(2024); recently, several potential altermagnetic materials have been identified by ab initio simulations Ma et al.
+(2024); Fedchenko et al.
 (2024); Šmejkal et al.
-(2022a); Lee et al.
-(2024).
-A characteristic property of $d$-wave altermagnets is $C_{4}\mathcal{T}$ symmetry, which breaks $\mathcal{T}$ and four-fold rotational ($C_{4}$) symmetry individually but preserve the combination of them.
-The interplay between the altermagnets and superconductivity has been theoretically investigated Zhu et al.
+(2022c); Hodt and Linder (2024); Zhu et al.
 (2023); Beenakker and Vakhtel (2023); Zhang et al.
 (2024); Ouassou et al.
+(2023); Sun et al.
 (2023); Papaj (2023); Li and Liu (2023); Giil et al.
-(2024).
-For ALM/SC heterostructures, we find that a nonvanishing out-of-plane spin magnetization can be generated in the second-order response to the driving supercurrent.
-This is in sharp contrast to the case of polar or gyrotropic superconductors Edelstein (1995); He et al.
-(2019); He and Law (2020).
-This work is organized as follows: First we present a general theory of supercurrent-induced spin generation including both the linear (first order) and nonlinear (second order) terms.
-Then we focus on the case in a specific ALM/SC realization.
-Remarkably, due to $C_{4}\mathcal{T}$ symmetry in altermagnets, we find that the second-order response $\chi_{ab}^{z}$ provides the leading order contribution to the out-of-plane spin magnetization; both equilibrium out-of-plane magnetization and linear-order out-of-plane magnetoelectric effects vanish.
-Additionally, due to the presence of a weak SOI in real altermagnets, the linear magnetoelectric effect also shows up in the in-plane direction.
-Furthermore, in the weak oscillating regime, driving an AC supercurrent can induce the second harmonic and rectified magnetization.
-Finally, we discuss possible material candidates that include thin films of the altermagnet RuO2 and KRu4O8 on an $s$-wave superconductor.
-We propose that ALM/SC devices allow for the non-dissipative electric control of magnetization, which is important for high density magnetic memories and offers tantalizing possibilities for spin transport in which Joule heating and dissipation are minimized.
-Superconducting magnetoelectric effect—We first examine the superconducting magnetoelectric effect within an effective Bogoliubov-de Gennes (BdG) framework.
-We note that our treatment is general and is agnostic to the precise microscopics of the superconducting gap: the superconducting state can be intrinsic or proximity induced.
-In what follows, we describe either case phenomenologically via the pairing potential $\Delta_{\bm{k}}$.
-The finite-$\bm{q}$ BdG Hamiltonian in the Nambu basis $(\hat{c}_{\bm{k}+\bm{q}/2,\uparrow},\hat{c}_{\bm{k}+\bm{q}/2,\downarrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\uparrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\downarrow})$ reads
+(2024); Sukhachov et al.
+(2024); Giil and Linder (2024); Hodt et al.
+(2024); Zyuzin (2024) characterized by a large spin-splitting of order hundreds of millielectronvolts Šmejkal et al.
+(2022a).
+As we describe below, the large spin-splitting in altermagnets naturally produces a pronounced second-order NSM $\chi_{ab}^{z}$ that dominates the out-of-plane spin magnetization of ALM/SC heterostructures.
+Indeed, when its $d$-wave $C_{4}\mathcal{T}$ symmetry is preserved, we find $\chi_{ab}^{z}$ provides the leading order contribution to the out-of-plane spin magnetization sharply contrasting with polar or gyrotropic superconductors Edelstein (1995); He et al.
+(2019); He and Law (2020) where the linear magnetoelectric effect dominates.
+Candidate altermagnetic materials include RuO2 and KRu4O8 thin films that can produce sizeable NSM susceptibilities.
+General theory of nonlinear superconducting magnetoelectric effect.—We first examine the superconducting magnetoelectric effect within an effective Bogoliubov-de Gennes (BdG) framework.
+We note that our treatment is general and is agnostic to the precise microscopics of the superconducting states, which can be intrinsic or proximity induced.
+In what follows, we describe either case phenomenologically via the pairing potential $\hat{\Delta}_{\bm{k}}$.
+The finite-$\bm{q}$ BdG Hamiltonian in the Nambu basis $\Psi_{\bm{k},\bm{q}}=(\hat{c}_{\bm{k}+\bm{q}/2,\uparrow},\hat{c}_{\bm{k}+\bm{q}/2,\downarrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\uparrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\downarrow})^{T}$ reads
 $\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}=\left(\begin{matrix}{}H_{\bm{k}+\bm{q}/2}&\hat{\Delta}_{\bm{k}}\\ \hat{\Delta}^{\dagger}_{\bm{k}}&-H^{*}_{-\bm{k}+\bm{q}/2}\end{matrix}\right),$ (2)
-where $H_{\bm{k}}$ is the Bloch Hamiltonian of the normal state, and $\bm{q}$ is the momentum of the Cooper pair.
-As we will see below, altermagnetic $H_{\bm{k}}$ enables the NSM effect.
+where $H_{\bm{k}}$ is the Bloch Hamiltonian in the normal state, and $\bm{q}$ is the momentum of the Cooper pair.
 For simplicity we have used a $\bm{q}$-independent gap function $\hat{\Delta}_{\bm{k}}$, which is valid for weak values of the applied current Chirolli et al.
 (2022).
-Table 1: Symmetry restrictions for three components of $\bm{M}$ in a two-dimensional system.
-Note that $C_{4}\mathcal{T}$ corresponds to the case for altermagnets that we focus on in our work.
-1(2) indicates first (second) order response.
-$\checkmark$ indicates allowed and $\times$ indicates forbidden.
-$\mathcal{P}$ $\mathcal{T}$ $\mathcal{PT}$ $C_{4}\mathcal{T}$
-$M_{x}$(2) ✓(1) $\times$ ✓(1)
-$M_{y}$(2) ✓(1) $\times$ ✓(1)
-$M_{z}$(2) ✓(1) $\times$ ✓(2)
-The total free energy $\mathcal{F}$ of the ALM/SC heterostructure comprises two parts: the superconductor component $\mathcal{F}_{\mathrm{sc}}$ and the material (i.e.
-altermagnet) component $\mathcal{F}_{\mathrm{al}}$ so that $\mathcal{F}=\mathcal{F}_{\mathrm{sc}}+\mathcal{F}_{\mathrm{al}}$.
+The free energy $\mathcal{F}$ of the heterostructure [Fig.
+1(a)] comprises two parts: the superconductor component $\mathcal{F}_{\mathrm{sc}}$ and the material (i.e.
+altermagnet) component $\mathcal{F}_{m}$ so that $\mathcal{F}=\mathcal{F}_{\mathrm{sc}}+\mathcal{F}_{m}$.
 The supercurrent can be tracked as $\bm{J}_{s}=\nu_{s}\bm{q}$ with the condensate superfluid stiffness $\nu_{s}=\partial^{2}\mathcal{F}_{\mathrm{sc}}(\bm{q})/\partial^{2}\bm{q}$.
 As a result, observables that scale with $q$ below are induced by the supercurrent at linear order while $q^{2}$ indicates a second-order nonlinear response.
-The magnetic response of the heterostructure can be obtained in the standard fashion by analyzing the set of parametric partition function $\mathcal{Z}(\bm{h})$
+To obtain the magnetic response, we analyze the set of parametric partition function $\mathcal{Z}(\bm{h})$
 $\mathcal{Z}(\bm{h})=e^{-\beta\{\mathcal{F}_{\mathrm{sc}}-\frac{1}{2\beta}\sum_{\bm{k},\bm{q},n}\mathrm{Tr}[\mathrm{log}G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})]\}},$ (3)
 where $\bm{h}$ is a parameter describing an auxiliary Zeeman field with a corresponding Zeeman energy $\Sigma(\bm{h})=\mathrm{diag}(g_{s}\mu_{B}\bm{h}\cdot\hat{s},-g_{s}\mu_{B}\bm{h}\cdot\hat{s}^{*})/2$.
-Here $g_{s}=2$ is the Landé $g$ factor, $\mu_{B}$ being the Bohr magneton, and the Gor’kov’s Green’s functions are $G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})=\mathcal{G}^{-1}(\bm{k},\bm{q},i\omega_{n})+\Sigma(\bm{h})$ and $\mathcal{G}(\bm{k},\bm{q},i\omega_{n})=(i\omega_{n}-\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}})^{-1}$ with the Matsubara frequency $\omega_{n}=(2n+1)\pi/\beta$.
+Here $g_{s}=2$ is the Landé $g$ factor, $\mu_{B}$ is the Bohr magneton, and the Gor’kov’s Green’s functions are $G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})=\mathcal{G}^{-1}(\bm{k},\bm{q},i\omega_{n})+\Sigma(\bm{h})$ and $\mathcal{G}(\bm{k},\bm{q},i\omega_{n})=(i\omega_{n}-\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}})^{-1}$ with the Matsubara frequency $\omega_{n}=(2n+1)\pi/\beta$.
 $\beta=1/k_{B}T$ with $T$ being the temperature.
@@
 Here $\eta_{a}=\mathrm{diag}(s_{a},-s_{a}^{*})$ are the generalized spin Pauli matrices in the Nambu space.
-To extract the linear and nonlinear responses systematically, we expand the BdG Hamiltonian as
-$\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}=\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}=0}+\frac{1}{2}q_{a}\hat{v}_{a}+\frac{1}{8}q_{a}q_{b}\hat{w}_{ab},$ (5)
-where $\hat{v}_{a}=\mathrm{diag}[\mathcal{V}_{a}(\bm{k}),-\mathcal{V}_{a}^{*}(-\bm{k})]$ and $\hat{w}_{ab}=\mathrm{diag}[\mathcal{W}_{ab}(\bm{k}),-\mathcal{W}_{ab}^{*}(-\bm{k})]$ with $\mathcal{V}_{a}(\bm{k})=\partial H_{\bm{k}}/\partial k_{a}$ and $\mathcal{W}_{ab}(\bm{k})=\partial^{2}H_{\bm{k}}/\partial k_{a}\partial k_{b}$.
-At linear order in $\bm{q}$, we find $\delta M_{a}^{(1)}=\alpha_{ab}q_{b}$, where the first-order spin susceptibility $\alpha_{ab}$ reads
+We then extract the linear and nonlinear responses systematically, and at linear order in $\bm{q}$ we find $\delta M_{a}^{(1)}=\alpha_{ab}q_{b}$, where the first-order spin susceptibility $\alpha_{ab}$ reads
 $\alpha_{ab}=-\frac{g_{s}\mu_{B}}{8\beta}\sum_{n\bm{k}}\mathrm{Tr}[\eta_{a}\mathcal{G}_{0}\hat{v}_{b}\mathcal{G}_{0}],$ (5)
 where $\mathcal{G}_{0}\equiv\mathcal{G}(\bm{k},0,i\omega_{n})=(i\omega_{n}-\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}=0})^{-1}$ for brevity.
+Here $\hat{v}_{a}=\mathrm{diag}[\mathcal{V}_{a}(\bm{k}),-\mathcal{V}_{a}^{*}(-\bm{k})]$ with $\mathcal{V}_{a}(\bm{k})=\partial H_{\bm{k}}/\partial k_{a}$.
 This formula is consistent with the results in Ref.
@@
 (1) with the second-order spin susceptibility $\chi$ as
-$\chi_{ab}^{c}=-\frac{g_{s}\mu_{B}}{32\beta}\!\sum_{n\bm{k}}\!\mathrm{Tr}[\eta_{c}(\mathcal{G}_{0}\hat{w}_{ab}\mathcal{G}_{0}+2\mathcal{G}_{0}\hat{v}_{a}\mathcal{G}_{0}\hat{v}_{b}\mathcal{G}_{0})].$ (7)
-We note that that $\chi_{ab}^{c}$ in Eq.
-(7) is general for arbitrary model Bloch Hamiltonians as well as pairing potentials.
-In particular, Eq.
-(7) applies for both conventional pairing as well as unconventional pairing $\hat{\Delta}_{\bm{k}}$.
-However, in the next section we will take a simple $s$-wave pairing $\hat{\Delta}_{\bm{k}}=\Delta_{0}is_{y}$ with a $\bm{k}$-independent pairing gap as an illustrative example to emphasize the effect of the magnetic order.
-Detailed derivations can be found in Supplementary Material Not .
+$\chi_{ab}^{c}=-\frac{g_{s}\mu_{B}}{32\beta}\!\sum_{n\bm{k}}\!\mathrm{Tr}[\eta_{c}(\mathcal{G}_{0}\hat{w}_{ab}\mathcal{G}_{0}+2\mathcal{G}_{0}\hat{v}_{a}\mathcal{G}_{0}\hat{v}_{b}\mathcal{G}_{0})],$ (6)
+where $\hat{w}_{ab}=\mathrm{diag}[\mathcal{W}_{ab}(\bm{k}),-\mathcal{W}_{ab}^{*}(-\bm{k})]$ with $\mathcal{W}_{ab}(\bm{k})=\partial^{2}H_{\bm{k}}/\partial k_{a}\partial k_{b}$.
+We note that $\chi_{ab}^{c}$ in Eq.
+(6) is general for various model Bloch Hamiltonians and applies for both conventional pairing as well as unconventional pairing $\hat{\Delta}_{\bm{k}}$.
+In the following we will take a simple $s$-wave pairing $\hat{\Delta}_{\bm{k}}=\Delta_{0}is_{y}$ with a $\bm{k}$-independent pairing gap as an illustrative example to emphasize the effect of the magnetic order.
+More details of the derivations can be found in Supplementary Material (SM) Not .
+Table 1: Symmetry restrictions for three components of $\bm{M}$ in a two-dimensional system.
+Note that $C_{4}\mathcal{T}$ corresponds to the case for altermagnets that we focus on in our work.
+Centrosymmetric $D_{3d}$ group corresponds to the kagome lattice.
+1(2) indicates first (second) order response.
+$\checkmark$ indicates allowed and $\times$ indicates forbidden.
+$\mathcal{P}$ $\mathcal{T}$ $\mathcal{PT}$ $C_{4}\mathcal{T}$ $D_{3d}$
+$M_{x}$(2) ✓(1) $\times$ ✓(1) ✓(2)
+$M_{y}$(2) ✓(1) $\times$ ✓(1) ✓(2)
+$M_{z}$(2) ✓(1) $\times$ ✓(2) $\times$
 It is instructive to examine the symmetry requirements for both the linear $\alpha_{ab}$ and second-order NSM $\chi_{ab}^{c}$ summarized in Table 1.
@@
 (5).
-Importantly, however, the Rashba SOI induced linear magnetoelectric effect is purely in-plane ($\alpha_{zx}=\alpha_{zy}=0$).
-This means that for out-of-plane magnetatization responses, the second order NSM $\chi_{ab}^{z}$ is the leading order term in the altermagnetic systems.
-As we will see below, it dominates the out-of-plane magnetization responses.
-Figure 2: (a) The fully gapped BdG energy spectrum of ALM/SC from Eq.
+Importantly, the Rashba SOI induced linear magnetoelectric effect is purely in-plane ($\alpha_{zx}=\alpha_{zy}=0$).
+This means that for out-of-plane magnetic responses, the second order NSM $\chi_{ab}^{z}$ is the leading order term in the altermagnetic systems.
+As we will see below, it dominates the out-of-plane magnetic responses.
+We note that Ref.
+Zyuzin (2024) recently discussed how $d$-wave magnetic order in a superconductor induces a second-order NSM response.
+Our general formulation reveals that such NSM responses also extend beyond altermagnets.
+An illustrative example involves materials in the $D_{3d}$ group; in such systems, we find that NSM persists and dominates the in-plane magnetic responses, see below.
+Figure 2: (a) The BdG energy spectrum of ALM/SC from Eq.
 (2) with Bloch Hamilotnian in Eq.
-(8) at $k_{y}=0$ using a phenomelogical gap $\hat{\Delta}_{\bm{k}}=\Delta_{0}is_{y}$.
+(7) at $k_{y}=0$ with spin-singlet pairing $\Delta_{0}$.
 (b) The spin magnetization $M_{z}$ as a function of $q_{x}$ from Eq.
@@
 The temperature is set to be $T=0.3T_{c}$.
-NSM in ALM/SC heterostructure—In the set-up illustrated in Fig.
-1(a), an altermagnet thin film is in contact with a conventional $s$-wave superconductor.
+In our calculations, we adopt the BCS temperature dependence of $\Delta_{0}$ with $\Delta(T)=\Delta_{0}\tanh(1.74\sqrt{T_{c}/T-1})$ and $\Delta_{0}=1.76k_{B}T_{c}$.
+NSM in ALM/SC heterostructure.—To exemplify NSM, we first focus on the ALM/SC set-up illustrated in Fig.
+1(a): an altermagnet thin film is in contact with a conventional $s$-wave superconductor.
 To characterize the physics of quasi-2D planar $d_{x^{2}-y^{2}}$-wave altermagnet, we adopt a generic two-band Hamiltonian Šmejkal et al.
 (2022a), yielding
-$H_{\bm{k}}=[ta^{2}(k_{x}^{2}+k_{y}^{2})-\mu]s_{0}+J_{\mathrm{ex}}a^{2}(k_{x}^{2}-k_{y}^{2})s_{z}.$ (8)
+$H_{\bm{k}}=(ta^{2}\bm{k}^{2}-\mu)s_{0}+J_{\mathrm{ex}}a^{2}(k_{x}^{2}-k_{y}^{2})s_{z}.$ (7)
 Here $t$ parameterizes the usual kinetic energy, $J_{\mathrm{ex}}$ denotes the $d_{x^{2}-y^{2}}$ exchange magnetic order and $\mu$ is the Fermi energy.
@@
 (2024).
-Here, $s$ matrices operate on the spin degree of freedom.
 This Hamiltonian breaks time reversal symmetry $\mathcal{T}=-is_{y}\mathcal{K}$ ($\mathcal{K}$ is the complex conjugate) and four-fold rotation symmetry $C_{4}=e^{i\pi s_{z}/4}$, but preserves the combination of them as $C_{4}\mathcal{T}$.
@@
 2(a).
-Although the pairing potential is purely $s$-wave, the induced superconducting correlations can be obtained as $F=[\psi_{0}+\bm{d}(\bm{k})\cdot\bm{s}]is_{y}$, where $\psi_{0}$ and $\bm{d}$ vector parameterize the spin-singlet and spin-triplet pairing correlations, respectively Gor’kov and Rashba (2001); Frigeri et al.
+Although the pairing potential is purely $s$-wave, the induced superconducting correlations can be obtained as $F(\bm{k},\omega)=[\psi_{0}+i\omega\bm{d}(\bm{k})\cdot\bm{s}]is_{y}$, where $\psi_{0}$ and $\bm{d}$ vector parameterize the spin-singlet and spin-triplet pairing correlations, respectively Bergeret et al.
+(2005); Gor’kov and Rashba (2001); Frigeri et al.
 (2004); Zhou et al.
-(2016).
-We have $d=(0,0,d_{z})$ with $d_{z}\propto J_{\mathrm{ex}}(k_{x}^{2}-k_{y}^{2})$, indicating that the $d$-wave magnetic order gives rise to the spin-triplet pairing correlation in altermagnets due to the proximity effect.
+(2016); Linder and Balatsky (2019).
+We have $d=(0,0,d_{z})$ with $d_{z}\propto J_{\mathrm{ex}}(k_{x}^{2}-k_{y}^{2})$, indicating that the $d$-wave magnetic order gives rise to the odd-frequency spin-triplet pairing correlation in altermagnets due to the proximity effect.
 To illustrate NSM effect in the ALM/SC heterostructure, we numerically study the supercurrent induced magnetization in Fig.
@@
 (7).
-By applying the current along $x$ direction as $\bm{q}=(q_{x},0)$, we calculate the magnetization directly from Eq.
-(4) at $J_{\mathrm{ex}}=0.2t$ as shown in Fig.
+By applying the current along $x$ direction as $\bm{q}=(q_{x},0)$, we plot the magnetization directly from Eq.
+(4) at $J_{\mathrm{ex}}=0.2t$ in Fig.
 2(b).
-One finds that $M_{z}\sim q_{x}^{2}$ as expected from Fig.
+We find that $M_{z}\sim q_{x}^{2}$ as expected from Fig.
 1(b).
@@
 (4)].
-In Fig.
-2(c), we show the second-order spin susceptibility $\chi_{xx}^{z}$ as a function of $J_{\mathrm{ex}}$ from Eq.
-(7).
-Generally, $\chi_{xx}^{z}$ starts to grow linearly with $J_{\mathrm{ex}}$ as shown in the zoom-in inset of Fig.
-2(c).
+[…]
+$\begin{split}f(\zeta)=&\frac{\beta^{3}}{64\pi}\zeta\mathrm{sech}^{2}\gamma\Big{[}\frac{4}{\beta^{2}}+t^{2}\zeta^{4}(2-3\mathrm{sech}^{2}\gamma)\\ &+\frac{4t\zeta^{2}(\mu-t\zeta^{2})\tanh\gamma}{\gamma}\Big{]},\end{split}$ (8)
+where $\gamma=\beta\sqrt{\Delta(T)^{2}+(t\zeta^{2}-\mu)^{2}}/2$.
 Furthermore, we also calculate the first-order spin susceptibility $\alpha_{xy}$ (purple) which is shown in Fig.
@@
 As expected, $\alpha_{xy}$ directly depends on $\lambda$; for small $\lambda$, $\alpha_{xy}$ increases linearly with $\lambda$ and vanishes when $\lambda=0$.
-In contrast, for NSM $\chi_{xx}^{z}$ is finite for $\lambda=0$ and exhibits a weak dependence on $\lambda$; $\chi_{xx}^{z}$ becomes smaller when $\lambda$ is nonzero since Rashba SOI tends to pin the spin in the in-plane direction and weakens the out-of-plane spin polarization.
-In our calculations, we adopt the BCS temperature dependence (for the $s$-wave superconductor substrate) of $\Delta_{0}$ with $\Delta(T)=\Delta_{0}\tanh(1.74\sqrt{T_{c}/T-1})$ and $\Delta_{0}=1.76k_{B}T_{c}$.
-Figure 3: The second-order spin susceptibility $\chi_{xx}^{z}$ as a function of the temperature $T$ at $(J_{\mathrm{ex}},\Delta_{0},\mu)=(0.02,0.1,0.3)t$.
-The purple diamonds denotes the numerical results from Eq.
-(7) and the orange line denotes the analytical result from Eq.
+The role of $\lambda$ (and more generally other SOI) for NSM $\chi_{xx}^{z}$ is more nuanced.
+At zero temperature, $\chi_{xx}^{z}$ vanishes as $\chi_{xx}^{z}\sim\beta^{3}\mathrm{sech}^{2}\beta$ [Eq.
+(8)] when $\lambda=0$ since $[\eta_{z},\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}]=0$ rendering $\langle\eta_{z}\rangle$ conserved.
+When $\lambda\neq 0$ is turned on, $\langle\eta_{z}\rangle$ is no longer conserved turning on $\chi_{xx}^{z}$.
+In contrast, at finite temperature $\chi_{xx}^{z}$ is sizeable even for $\lambda=0$ yielding a nonzero $M_{z}$ that arises from thermal fluctuations at finite temperature in the current-carrying state.
+Indeed, $\chi_{xx}^{z}$ at finite temperature (e.g., $T=0.3T_{\rm c}$) in Fig.
+2(d) exhibits a weak dependence on $\lambda$.
+This evidences the large altermagentic spin splitting Šmejkal et al.
+(2022a) of order several hundred millielectronvolts captured by $J_{\rm ex}\gg\lambda$ dominates the magnetic order yielding a robust NSM.
+NSM in a noncollinear antiferromagnet.—So far we have focused on NSM in $d$-wave altermagnets—a type of collinear antiferromagnet.
+The general NSM in Eq.
+(6) can exist in a larger class of magnet/superconductor heterostructures (see e.g., Table 1 as well as SM Not for a detailed discussion).
+Indeed, based on Table 1, we find that for centrosymmetric noncollinear antiferromagnet with $D_{3d}$ symmetry, an in-plane NSM with $\chi_{xx}^{y}=\chi_{xy}^{x}=\chi_{yx}^{x}=-\chi_{yy}^{y}\neq 0$ can manifest.
+To capture the physics of a metal with noncollinear magnetic order on a kagome lattice, we adopt the low energy Hamiltonian Lee et al.
+(2024a); Not as
+$H_{\bm{k}}=(ta^{2}\bm{k}^{2}-\mu)s_{0}+ma^{2}[2k_{x}k_{y}s_{x}+(k_{x}^{2}-k_{y}^{2})s_{y}],$ (9)
+where $m$ captures the noncollinear magnetic order.
+Note that the noncollinear magnetic order induced winding spin texture on the Fermi surface [Fig.
+3(a)] sharply contrasts with the collinear magnetic order in $d$-wave altermagnets [Fig.
+1(c)].
+As we now explain, this difference in magnetic order leads to an in-plane NSM.
+To see this, we follow the same procedure as described above and utilize the parameters $(m,\mu,\Delta_{0})=(0.2,0.3,0.1)t$.
+We obtain an in-plane NSM in Fig.
+3(b) that dominates the magnetoelectricity; indeed, as expected from Table.
+1 (last column), $M_{y}$ is quadratic in $q_{x}$ – it is the leading order in-plane magnetoelectric term.
+In-plane NSM is directly controlled by the noncollinear magnetic order as evidenced in Fig.
+3(c), where $\chi_{xx}^{y}$ turns on with $m$.
+As a result, we find that NSM arises in the presence of a variety of magnetic orders: such as $J_{\mathrm{ex}}$ (altermagnet, collinear) or $m$ (kagome, noncollinear) yielding both out-of-plane as well as in-plane NSM responses.
+Figure 3: (a) The Fermi surface associated with the winding spin texture of Eq.
 (9).
-At low temperatures, $\chi^{z}_{xx}\sim\beta^{3}\mathrm{sech}^{2}\beta$.
-[…]
-$\begin{split}f(x)=&\frac{\beta^{3}J_{\mathrm{ex}}}{64\pi}x\mathrm{sech}^{2}\gamma\Big{[}\frac{4}{\beta^{2}}+t^{2}x^{4}(2-3\mathrm{sech}^{2}\gamma)\\ &+\frac{4tx^{2}(\mu-tx^{2})\tanh\gamma}{\gamma}\Big{]},\end{split}$ (9)
-where $\gamma=\beta\sqrt{\Delta(T)^{2}+(tx^{2}-\mu)^{2}}/2$.
-By employing Eq.
-(9), we plot the $\chi_{xx}^{z}$ as a function of temperature in Fig.
-3 (orange solid curve) with $J_{\mathrm{ex}}=0.02t$.
-By way of comparison, we also show a fully numerical plot of Eq.
-(7) (purple diamonds); both agree with each other.
-Interestingly, $\chi_{xx}^{z}$ vanishes at zero temperature.
-At low temperatures ($\beta\Delta_{0}\gg 1$), $\chi_{xx}^{z}$ starts to grows with $T$ scaling as $\chi_{xx}\sim\beta^{3}\mathrm{sech}^{2}\beta$.
-However, at $T/T_{c}\sim 0.6$, it reaches a peak after which it rapidly diminishes with temperature.
-This non-monotonic behavior contrasts sharply with that of the linear magnetoelectric effect in Rashba-type superconductors He et al.
-(2019); Not that instead saturates at low temperature.
-In superconductors, supercurrent flow can arise due to either an external applied magnetic field or an explicit transport current Anthore et al.
-(2003).
-In the former case, there is a screening supercurrent on the surface of the superconductor.
-In the latter case, the global excitation of the superconductor gives a uniform Cooper pair momentum $\bm{q}$.
-Specifically, for slowly varying AC supercurrent as $\bm{J}_{s}(t)=\bm{J}_{\omega}\cos\omega t$ with $\omega\ll\Delta_{0}$, we expect an oscillatory Cooper pair momentum $\bm{q}(t)=\bm{q}_{\omega}\cos\omega t$.
-This will allow to obtain a second harmonic magnetization in the weak oscillating regime with $\chi_{ab}^{c}(2\omega)=\chi_{ab}^{c}/4$.
-Candidate materials.—We anticipate that NSM we discuss here can be realized in ALM/SC formed out of readily available $d$-wave altermagnetic materials Šmejkal et al.
-(2022a).
-As an example, consider the altermagnet RuO2 with in-plane lattice constant $a=4.5\AA$, and $t=2.5$ eV, $J_{\mathrm{ex}}=0.5$ eV Šmejkal et al.
+Blue (yellow) color denotes the upper (lower) band.
+(b) The spin magnetization $M_{y}$ as a function of $q_{x}$ from Eq.
+(4) (solid orange) aligns with $\chi^{y}_{xx}q_{x}^{2}$ from Eq.
+(6) (dashed green).
+(c) $\chi^{y}_{xx}$ as a function of noncollinear magnetic order $m$.
+We set $(\Delta_{0},\mu)=(0.1,0.3)t$ and $T=0.3T_{c}$.
+Discussion.—We expect that NSM can be realized in several concrete materials.
+As an example, considering the altermagnet RuO2 thin film with lattice constant $a=4.5\AA$, $c=3.1\AA$, and $t=2.5$ eV, $J_{\mathrm{ex}}=0.5$ eV Šmejkal et al.
 (2022b); Zhang et al.
 (2024).
-Using a pairing gap of $\Delta_{0}=1$ meV, $\mu=10$ meV, $\lambda=10$ meV and $T=0.3T_{c}$ we estimate $\chi_{xx}^{z}\approx 0.07\mu_{B}$.
+Using a pairing gap of $\Delta_{0}=1$ meV, $\mu=0.1$ eV, $\lambda=40$ meV and $T=0.3T_{c}$ we estimate $\chi_{xx}^{z}\approx 0.1\mu_{B}$.
 For a Cooper pair momentum $q\approx 10^{-3}\AA^{-1}$ Zhu et al.
-(2021), the induced out-of-plane spin density in the RuO2 thin film/superconductor heterostructure is $\sim 0.7\times 10^{-5}\mu_{B}/\mathrm{nm}^{2}$.
+(2021), the induced out-of-plane spin density in the RuO2 thin film/superconductor heterostructure is $\sim 0.3\times 10^{-4}\mu_{B}/\mathrm{nm}^{3}$.
 Note that this value is comparable to the spin generation in MnBi2Te4 and other noncentrosymmetric ferromagnetic systems Xiao et al.
@@
 (2011).
-Another example is KRu4O8 with the parameter values: $t=0.05$ eV, $J_{\mathrm{ex}}=0.018$ eV, and $a=9.9\AA$.
+Another example is KRu4O8 with the parameter values: $t=0.05$ eV, $J_{\mathrm{ex}}=0.018$ eV, and $a=9.9\AA$ Zhang et al.
+(2024).
 We find $\chi_{xx}^{z}\approx 0.01\mu_{B}$ and out-of-plane spin density similar to that discussed above.
-These magnitudes are estimated qualitatively using a low-energy effective description of altermagnets in the small $J$ and weak doping (small $\mu$) regime as well as a gapped BdG spectrum.
-In real altermagnets with large $J$ and high doping, the BdG spectrum becomes gapless.
-We anticipate this may enhance NSM with additional contribution from the quasiparticles excited at finite temperature.
-In this work, we have proposed a novel nonlinear superconducting magnetoelectric effect.
-Naturally occurring in ALM/SC heterostructures, this second-order spin generation in response to the supercurrent becomes the leading order contribution to the magnetization along the out-of-plane direction.
-Supercurrent induced magnetization can be readily detected by using a superconducting quantum interference device (SQUID) perpendicular to the altermagnet surface, which can be employed to probe the magnetic flux change Chirolli et al.
+The induced NSM magnetization densities (of order $\sim 10^{-4}\mu_{B}/\mathrm{nm}^{3}$ corresponding to equivalent magnetic fields $\sim 1\mu{\rm T}$) can be readily detected by superconducting quantum interference devices (SQUID) Chirolli et al.
+(2022); Vasyukov et al.
+(2013) or diamond nitrogen-vacancy center-based magnetometers Balasubramanian et al.
 (2009).
-Interestingly, nonlinear spin generation that is second order in an applied electric field can be realized in normal-state (non-superconducting) altermagnets with $C_{4}\mathcal{T}$ symmetry Xiao et al.
+Naturally, high quality interfaces should be used, since disorder can locally disrupt the superconducting pairing and suppress the supercurrent.
+Additionally, in our analysis above, we have focussed on ideal ALM/SC heterostructures; in realistic systems, surface reconstruction may ensue.
+For example, reconstruction can produce to interfacial spin-orbit coupling that breaks local mirror symmetry Linder and Yokoyama (2011), inducing additional linear magnetoelectric responses including in the out-of-plane direction (see SM Not of interface effects).
+Nevertheless, we expect the large altermagnetic spin-splitting Šmejkal et al.
+(2022a) readily dominates the response rendering NSM robust.
+Even in the presence of an out-of-plane linear magnetoelectric response, we anticipate that NSM can be distinguished by dynamical probes: e.g., oscillatory Cooper pair momentum $\bm{q}(t)=\bm{q}\cos\omega t$ for slowly varying AC supercurrent with $\omega\ll\Delta_{0}$.
+In this weak oscillating regime, a rectified as well as a second harmonic magnetization appear: $\delta M_{c}^{(2)}(t)=\chi^{c}_{ab}q_{a}q_{b}(1+\cos 2\omega t)/2$.
+Both rectified as well as second harmonic depart from that of the conventional linear magnetoelectric effect, which produces an oscillating magnetization at frequency $\omega$.
+This distinction is important practically since it enables to directly identify NSM even in systems where NSM is not the leading order magnetoelectric effect: e.g., conventional linear magnetoelectric effect produces a magnetization that averages to zero over one cycle, while the rectified NSM does not vanish allowing to directly identify its contribution; similarly locking-in to $2\omega$ allows to directly distinguish between linear and nonlinear superconducting magnetoelectric effects.
+While NSM is naturally realized in ALM/SC heterostructures (driven by its large spin-splitting), we find that it persists more generally for a wide range of other materials and symmetries (see SM Not ).
+Strikingly, when NSM is combined with the linear magnetoelectric effect, current induced magnetization can become non-reciprocal producing distinct magnetizations for $+\bm{q}$ and $-\bm{q}$ Not , a new tool for non-dissipative spintronics.
+Perhaps most interesting, however, is the NSM’s ability to distinguish dilineate magnetic orders: as a result, we expect NSM provides a promising approach for exploring the rich interplay between magnetism and superconductivity in unconventional magnetic materials that include kagome materials Yin et al.
+(2022), as well as materials in different spin space groups Xiao et al.
+(2024); Jiang et al.
+(2024); Chen et al.
 (2024).
-Our work shows that nonlinear spin generation survives in the superconducting state and is driven by a dissipationless supercurrent.
-Acknowledgements.—We thank Bo Yang and Wen-Yu He for helpful discussions.
+Acknowledgements.—We thank Bo Yang, Wen-Yu He, Naizhou Wang and Benjamin T.
+Zhou for helpful discussions.
 This work was supported by the Ministry of Education Singapore under its Academic Research Fund Tier 2 Grant No.
 MOE-T2EP50222-0011.
-Note added.—Recently, we became aware of an independent pre-print posted on the arXiv in Ref.
-Zyuzin (2024) which also found a second-order supercurrent-induced spin polarization in superconductors with $d$-wave magnetization.
 Supplementary Material for
@@
 1Division of Physics and Applied Physics, School of Physical and Mathematical Sciences, Nanyang Technological University, Singapore 637371
-I I. Theory of nonlinear superconducting Magnetoelectric effect
-I.1 A. Green’s function method
-In this section we provide more details for deriving the general formula of superconducting magnetoelectric effect, including both the linear and nonlinear terms.
+S1 Microscopic theory of superconducting magnetoelectric effect
+S1.1 Green’s function method
+We provide more details in deriving the general formula of superconducting magnetoelectric effect, including both the linear and nonlinear terms.
 The superconducting state can be intrinsic or proximity induced with a pairing function $\hat{\Delta}_{\bm{k}}$.
 $\hat{\Delta}_{\bm{k}}$ can be conventional $s$-wave with a uniform pairing gap $\hat{\Delta}_{\bm{k}}=\Delta_{0}is_{y}$ or unconventional pairing states.
-In the current-carrying state, the finite-$\bm{q}$ Bogoliubov-de Gennes (BdG) Hamiltonian in the Nambu basis $(\hat{c}_{\bm{k}+\bm{q}/2,\uparrow},\hat{c}_{\bm{k}+\bm{q}/2,\downarrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\uparrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\downarrow})$ reads
-$\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}=\left(\begin{matrix}{}H_{\bm{k}+\bm{q}/2}&\hat{\Delta}_{\bm{k}}\\ \hat{\Delta}^{\dagger}_{\bm{k}}&-H^{*}_{-\bm{k}+\bm{q}/2}\end{matrix}\right)$ (S1)
+In the current-carrying state, the finite-$\bm{q}$ mean field Hamiltonian is $H_{\mathrm{mf}}(\bm{q})=\frac{1}{2}\Psi_{\bm{k},\bm{q}}^{\dagger}\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}\Psi_{\bm{k},\bm{q}}$ in the Nambu basis $\Psi_{\bm{k},\bm{q}}=(\hat{c}_{\bm{k}+\bm{q}/2,\uparrow},\hat{c}_{\bm{k}+\bm{q}/2,\downarrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\uparrow},\hat{c}^{\dagger}_{-\bm{k}+\bm{q}/2,\downarrow})^{T}$.
+The Bogoliubov-de Gennes (BdG) Hamiltonian reads
+$\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}=\left(\begin{matrix}{}H_{\bm{k}+\bm{q}/2}&\hat{\Delta}_{\bm{k}}\\ \hat{\Delta}^{\dagger}_{\bm{k}}&-H^{*}_{-\bm{k}+\bm{q}/2}\end{matrix}\right),$ (S1)
 where $H_{\bm{k}}$ is the Bloch Hamiltonian of the normal state, $\bm{q}$ is the momentum of the Cooper pair, and $s_{y}$ is a Pauli matrix.
@@
 In the main text, we focus on the altermagnet/superconductor heterostructure.
-The total free energy $\mathcal{F}$ of the system contains the superconductor part $\mathcal{F}_{\mathrm{sc}}$ and the altermagnet part $\mathcal{F}_{\mathrm{al}}$ with $\mathcal{F}=\mathcal{F}_{\mathrm{sc}}+\mathcal{F}_{\mathrm{al}}$.
+Therefore the total free energy $\mathcal{F}$ of the system contains the superconductor part $\mathcal{F}_{\mathrm{sc}}$ and the material part $\mathcal{F}_{m}$ with $\mathcal{F}=\mathcal{F}_{\mathrm{sc}}+\mathcal{F}_{\mathrm{m}}$.
 The supercurrent $J_{s}=\nu_{s}\bm{q}$ with the superfluid density $\nu_{s}=\partial^{2}\mathcal{F}_{\mathrm{sc}}(\bm{q})/\partial^{2}\bm{q}$.
@@
 The magnetization $M$ can be obtained from the partition function as
-$\begin{split}M_{a}&=-\frac{1}{\beta}\frac{\partial}{\partial{h_{a}}}\mathrm{log}Z(\bm{h})\\ &=-\frac{1}{2\beta}\sum_{\bm{k},\bm{q},n}\frac{\partial}{\partial{h_{a}}}\mathrm{Tr}[\mathrm{log}G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})]\end{split}$ (S4)
+$\begin{split}M_{a}&=-\frac{1}{\beta}\frac{\partial}{\partial{h_{a}}}\mathrm{log}Z(\bm{h})\\ &=-\frac{1}{2\beta}\sum_{\bm{k},\bm{q},n}\frac{\partial}{\partial{h_{a}}}\mathrm{Tr}[\mathrm{log}G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})].\end{split}$ (S4)
 We decompose the Green’s function as $G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})=\mathcal{G}^{-1}(\bm{k},\bm{q},i\omega_{n})+\Sigma(\bm{h})$ with
-$\Sigma(\bm{h})=\left(\begin{array}[]{cc}\frac{1}{2}g_{s}\mu_{B}\bm{h}\cdot\bm{s}&0\\ 0&-\frac{1}{2}g_{s}\mu_{B}\bm{h}\cdot\bm{s}\end{array}\right),$ (S5)
+$\Sigma(\bm{h})=\left(\begin{array}[]{cc}\frac{1}{2}g_{s}\mu_{B}\bm{h}\cdot\bm{s}&0\\ 0&-\frac{1}{2}g_{s}\mu_{B}\bm{h}\cdot\bm{s}\end{array}\right).$ (S5)
 We can then expand the logarithm in the first order of $\bm{h}$ as $\mathrm{log}G^{-1}(\bm{k},\bm{q},\bm{h},i\omega_{n})=\mathrm{log}\mathcal{G}^{-1}(\bm{k},\bm{q},i\omega_{n})+\mathcal{G}(\bm{k},\bm{q},i\omega_{n})\Sigma(\bm{h})$.
@@
 We can then expand the BdG Hamiltonian as
-$\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}=\mathcal{H}_{\mathrm{BdG}}^{\bm{k}}+q_{a}\hat{v}_{a}/2+q_{a}q_{b}\hat{w}_{ab}/8+\mathcal{O}(\bm{q}^{3})$ (S7)
-The velocity operators are
+$\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}=\mathcal{H}_{\mathrm{BdG}}^{\bm{k}}+q_{a}\hat{v}_{a}/2+q_{a}q_{b}\hat{w}_{ab}/8+\mathcal{O}(\bm{q}^{3}),$ (S7)
+where the velocity operators are
 $\hat{v}_{a}=\left(\begin{array}[]{cc}\hat{\mathcal{V}}_{a}(\bm{k})&0\\ 0&-\hat{\mathcal{V}}^{*}_{a}(-\bm{k})\end{array}\right),\hat{w}_{ab}=\left(\begin{array}[]{cc}\mathcal{W}_{ab}(\bm{k})&0\\ 0&-\mathcal{W}_{ab}^{*}(-\bm{k})\end{array}\right).$ (S8)
@@
 $\chi_{ab}^{c}=-\frac{g_{s}\mu_{B}k_{B}T}{32}\!\sum_{n\bm{k}}\!\mathrm{Tr}[\eta_{c}(\mathcal{G}_{0}\hat{w}_{ab}\mathcal{G}_{0}+2\mathcal{G}_{0}\hat{v}_{a}\mathcal{G}_{0}\hat{v}_{b}\mathcal{G}_{0})].$ (S12)
-I.2 B. Analytical result
+S1.2 Analytical result
 In the previous subsection, we use the Green’s function method to give the general formulas of the linear and nonlinear magnetization.
-These formulas can be adopt to evaluate the spin susceptibilities for altermagnet/superconductor heterostructures.
+These formulas can be adopted to evaluate the spin susceptibilities for altermagnet/superconductor heterostructures.
 In this subsection, we provide the analytical solution of the nonlinear spin susceptibility $\chi^{z}_{ab}$.
-For the altermagnets, we can write down a general two-band model
-$H_{\bm{k}}=[ta^{2}(k_{x}^{2}+k_{y}^{2})-\mu]s_{0}+J_{ex}a^{2}[2k_{x}k_{y}\cos(2\theta)+(k_{x}^{2}-k_{y}^{2})\sin(2\theta)]s_{z}$ (S13)
-$t$ parameterizes the usual kinetic energy, $J_{ex}$ denotes the $d$-wave exchange magnetic order parameter and $\mu$ is the chemical potential.
+For ideal altermagnets, we can write down a general two-band model
+$H_{\bm{k}}=[ta^{2}(k_{x}^{2}+k_{y}^{2})-\mu]s_{0}+J_{ex}a^{2}[2k_{x}k_{y}\cos(2\theta)+(k_{x}^{2}-k_{y}^{2})\sin(2\theta)]s_{z},$ (S13)
+where $t$ parameterizes the usual kinetic energy, $J_{ex}$ denotes the $d$-wave exchange magnetic order parameter and $\mu$ is the chemical potential.
 $\theta$ is the angle between the altermagnetic orientation and the crystalline axes.
@@
 and
-$\hat{w}_{xx}=a^{2}\left(\begin{array}[]{cc}2ts_{0}+2J_{ex}\sin(2\theta)s_{z}&0\\ 0&-2ts_{0}-2J_{ex}\sin(2\theta)s_{z}\end{array}\right),\hat{w}_{xy}=a^{2}\left(\begin{array}[]{cc}2J_{ex}\cos(2\theta)s_{z}&0\\ 0&-2J_{ex}\cos(2\theta)s_{z}\end{array}\right),$ (S16)
+$\hat{w}_{xx}=a^{2}\left(\begin{array}[]{cc}2ts_{0}+2J_{ex}\sin(2\theta)s_{z}&0\\ 0&-2ts_{0}-2J_{ex}\sin(2\theta)s_{z}\end{array}\right),\hat{w}_{xy}=a^{2}\left(\begin{array}[]{cc}2J_{ex}\cos(2\theta)s_{z}&0\\ 0&-2J_{ex}\cos(2\theta)s_{z}\end{array}\right).$ (S16)
 […]
@@
 where $\gamma=\beta\sqrt{\Delta^{2}+(tx^{2}-\mu)^{2}}/2$.
-II II. Pairing correlations
+Figure S1: (a) The second-order spin susceptibility $\chi_{xx}^{z}$ as a function of the temperature $T$.
+The purple diamonds denotes the numerical results from Eq.
+(S12) and the orange line denotes the analytical result from Eq.
+(S17).
+At low temperatures, $\chi^{z}_{xx}\sim\beta^{3}\mathrm{sech}^{2}\beta$.
+(b) $\chi_{xx}^{z}$ as a function of the Fermi energy $\mu$ at $T=0.3,0.6T_{c}$.
+(c),(d) The calculations of $\chi_{xx}^{z}$ for $\lambda=0$ (c) and $\lambda=0.1t$ (d).
+In (c), $\chi_{xx}^{z}$ is suppressed as $T\rightarrow 0$ because of the factor $\frac{1}{T^{3}}\mathrm{sech}^{2}(\frac{1}{T})$.
+In (d), $\chi_{xx}^{z}$ saturate to a finite value as $T\rightarrow 0$ for nonzero SOI.
+Other Parameters: $J=0.02t$ for (a) and (b).
+$J=0.2t$ for (c) and (d).
+$\mu=0.3t,\Delta_{0}=0.1t$ for all four panels.
+By employing Eq.
+(S17), we plot $\chi_{xx}^{z}$ as a function of temperature in Fig.
+S1(a) (orange solid curve) with $J_{\mathrm{ex}}=0.02t$.
+By way of comparison, we plot a numerical evaluation of Eq.
+(S12) (purple diamonds); both agree with each other.
+At low temperatures ($\beta\Delta_{0}\gg 1$), $\chi_{xx}^{z}$ starts to grows with $T$ scaling as $\chi_{xx}\sim\beta^{3}\mathrm{sech}^{2}\beta$.
+Moreover, at $T/T_{c}\sim 0.6$, $\chi_{xx}^{z}$ reaches a peak after which it rapidly diminishes with temperature.
+This non-monotonic behavior contrasts sharply with that of the linear magnetoelectric effect in Rashba-type superconductors He et al.
+(2019) that instead saturates at low temperature (see discussion below).
+We also plot the $\mu$ dependence of $\chi_{xx}^{z}$ from Eq.
+(S17) in Fig.
+S1(b).
+Larger NSM response can be achieved by increasing the Fermi energy $\mu$ because of the enhanced density of states.
+From this analytical formula, we find that at low temperature, $\chi$ is suppressed as $\frac{1}{T^{3}}\mathrm{sech}^{2}(\frac{1}{T})$.
+This can be understood in the following way: for an ideal altermagnet ($\lambda=0$), $\eta_{z}$ is conserved (recall $\eta_{z}=\mathrm{diag}[s_{z},-s_{z}]$), namely $[\eta_{z},\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}]=0$, which indicates that nonzero $M_{z}$ arises from the thermal fluctuations in the current-carrying state at nonzero temperature.
+The suppression of $\chi_{xx}^{z}$ is shown in Fig.
+S1(c) with $\lambda=0$ (zero SOI).
+In the presence of SOI ($\lambda\neq 0$), $[\eta_{z},\mathcal{H}_{\mathrm{BdG}}^{\bm{k},\bm{q}}]\neq 0$, leading to finite $M_{z}$ even when $T\rightarrow 0$.
+This is shown in Fig.
+S1(d).
+S1.3 Effects from local mirror breaking
+When fabricating a hybrid magnetic material/superconductor heterostructure, surface reconstruction at the interface can ensue.
+For instance, surface reconstruction at the interface can break the mirror symmetry locally rendering the total system effectively noncentrosymmetric Linder and Yokoyama (2011).
+This yields an interfacial spin orbital interaction (SOI) $H_{\mathrm{SOI}}=V_{s}(\bm{k}\times\bm{\sigma})\cdot\hat{\bm{n}}$ where $\hat{\bm{n}}$ is an interface normal vector.
+When $\hat{\bm{n}}=\hat{z}$, $V_{s}$ describes the well-known Rashba-like SOI, which preserves $C_{4}\mathcal{T}$.
+When $\hat{\bm{n}}=\hat{x}$ or $\hat{y}$, $k_{y}\sigma_{z}$ or $k_{x}\sigma_{z}$ can emerge as additional terms which break $C_{4}\mathcal{T}$.
+Since we have studied the Rashba SOI in the main text and above, here we present a simple model to illustrate the effect from local mirror symmetry breaking.
+To be specific, we consider the following effective Bloch Hamiltonian
+$H_{\bm{k}}=[ta^{2}(k_{x}^{2}+k_{y}^{2})-\mu]s_{0}+\lambda a(k_{x}s_{y}-k_{y}s_{x})+J_{\mathrm{ex}}a^{2}(k_{x}^{2}-k_{y}^{2})s_{z}+(\gamma_{x}k_{x}+\gamma_{y}k_{y})a\sigma_{z},$ (S18)
+where $\gamma_{x}$ and $\gamma_{y}$ denote the mirror breaking terms.
+In Fig.
+S2(a), we plat $M_{z}$ with finite $\gamma_{x}$ and $\gamma_{y}=0$.
+It is clear that in the presence of $\gamma_{x}$, we have a general form for the current induced magnetization as $M_{z}=Aq_{x}^{2}+Bq_{x}+C$.
+The linear term parameterized by the coefficient B arises from nonzero $\gamma_{x}$ which breaks mirror symmetry $\mathcal{M}_{y}$.
+In Fig.
+S2(b), we plot $M_{z}$ with finite $\gamma_{y}$ and $\gamma_{x}=0$.
+The $\gamma_{y}$ term breaks mirror symmetry $\mathcal{M}_{x}$ but preserve $\mathcal{M}_{y}$, which forbids the linear term B.
+Note that in both cases, we have $M_{z}\neq 0$ at $q_{x}=0$, resulting from the fact that $C_{4}\mathcal{T}$ is nor preserved in the presence of $\gamma_{x}$ or $\gamma_{y}$.
+Importantly, in all cases, NSM is robust and can dominate the magnetoelectric response.
+Interestingly, the combination of both linear and nonlinear superconducting magnetoelectric effects in Fig.
+S2(a) leads to a non-reciprocal magneto-electric effect with magnetizations distinct for $+q$ and $-q$.
+Figure S2: NSM with local mirror symmetry breaking from interfacial spin-orbital interaction.
+In (a) $\gamma_{y}=0$ and in (b) $\gamma_{x}=0$.
+$\gamma_{x},\gamma_{y}$ denotes the strength of SOI for the mirror breaking.
+Parameters: $J=0.2t,\mu=0.3t,\lambda=0.05t,\Delta_{0}=0.1t$ and $T=0.3T_{c}$.
+Figure S3: The supercurrent induced magnetization for (a) Spin-singlet $s$-wave pairing with $\hat{\Delta}_{\bm{k}}=\Delta_{0}is_{y}$ at $\Delta_{0}=0.1t$, (b) Spin-singlet $s_{\pm}$-wave pairing with $\hat{\Delta}_{\bm{k}}=(\Delta_{0}+\Delta_{1}\bm{k}^{2})is_{y}$ at $\Delta_{0}=0.1t$ and $\Delta_{1}=0.05t$, (c) Spin-singlet $d$-wave pairing with $\hat{\Delta}_{\bm{k}}=\Delta_{0}(k_{x}^{2}-k_{y}^{2})is_{y}$ at $\Delta_{0}=0.1t$, (d) Spin-triplet chiral $p$-wave pairing with $\hat{\Delta}_{\bm{k}}=\Delta_{0}[(k_{x}+ik_{y})s_{z}]is_{y}$ at $\Delta_{0}=0.1t$.
+Other parameters: $(J_{\mathrm{ex}},\lambda,\mu)=(0.2,0.05,0.3)t$.
+The temperature is set to be $k_{B}T=0.2\Delta_{0}$ for all four panels.
+S1.4 NSM with unconventional pairing states
+In the main text, we focus on the nonlinear magnetic response with spin-singlet $s$-wave pairing.
+Here we present more examples for the NSM with different unconventional pairing states.
+We consider (1) Spin-singlet $s_{\pm}$-wave pairing with $\hat{\Delta}_{\bm{k}}=(\Delta_{0}+\Delta_{1}\bm{k}^{2})is_{y}$.
+(2) Spin-singlet $d$-wave pairing with $\hat{\Delta}_{\bm{k}}=\Delta_{0}(k_{x}^{2}-k_{y}^{2})is_{y}$.
+(3) Spin-triplet $p$-wave pairing with $\hat{\Delta}_{\bm{k}}=\Delta_{0}[(k_{x}+ik_{y})s_{z}]is_{y}$.
+For simplicity, we consider the cases that the pairing states do not break the $C_{4}\mathcal{T}$ symmetry of $H_{\mathrm{BdG}}$.
+In Fig.
+S3 we calculate the $M_{z}$ for these different pairing potentials.
+For these different superconducting pairing states, $M_{z}$ can be quantitatively different, but $M_{z}\sim q_{x}^{2}$ persists as a general behavior (quadratic in Cooper pair momentum).
+S1.5 Details of material calculation
+Here we provide more details of the two examples in the main text.
+For RuO2 with in-plane lattice constant $a=4.5\AA$, and $t=2.5$ eV, $J_{\mathrm{ex}}=0.5$ eV, we assume the Rashba SOI term is $\lambda=40$ meV, which is an estimation including both the intrinsic and interface contributions.
+Using a pairing gap of $\Delta_{0}=1$ meV and $T=0.3T_{c}$ we plot $\chi_{xx}^{z}$ as a function of Fermi energy $\mu$ in Fig.S4(a).
+Larger response for larger $\mu$ is found because of the enhanced density of states.
+For KRu4O8 we adopt the parameters: $t=0.05$ eV, $J_{\mathrm{ex}}=0.018$ eV, and $a=9.9\AA$.
+The SOI and pairing gap is the same as RuO2, and we also get a similar result in Fig.
+S4(b).
+Figure S4: The calculation of $\chi_{xx}^{z}$ [Eq.
+(S12)] for RuO2 (a) and KRu4O8 (b) as a function of Fermi energy $\mu$.
+For RuO2, $t=2.5$ eV, $J_{\mathrm{ex}}=0.5$ eV, $\lambda=40$ meV.
+For KRu4O8, $t=0.05$ eV, $J_{\mathrm{ex}}=0.018$ eV, $\lambda=40$ meV.
+We take $\Delta_{0}=1$meV and $T=0.3T_{c}$ for the two plots.
+We take RuO2 as an example to estimate the magnetization.
+For a Cooper pair momentum $q\approx 10^{-3}\AA^{-1}$, the induced out-of-plane spin density in the RuO2 thin film/superconductor heterostructure is $\sim 10^{-5}\mu_{B}/\mathrm{nm}^{2}$ at $\chi_{xx}^{z}=0.1\mu_{B}$.
+Considering a thin film with the lattice constant $c=0.31$nm, the magnetization in the standard unit is $\delta M=3\times 10^{-5}\mu_{B}/\mathrm{nm}^{3}$.
+Using this, we estimate the equivalent magnetic field as $\delta B=\mu_{0}\delta M\approx 0.4\times 10^{-6}$.
+S2 Proximity effect and pairing correlations
 Although we only consider the simplest conventional spin singlet pairing, i.e., $\Delta_{0}is_{y}$, here we show that the $d$-wave magnetic order will induce the spin-triplet correlations.
@@
 $\mathcal{G}(\bm{k},i\omega_{n})=(i\omega_{n}-H_{\mathrm{BdG}}^{\bm{k}})^{-1}=\begin{pmatrix}G_{e}(\bm{k},i\omega_{n})&F(\bm{k},i\omega_{n})\\ F^{\dagger}(\bm{k},i\omega_{n})&G_{h}(\bm{k},i\omega_{n})\end{pmatrix}.$ (S22)
-Substitute the BdG Hamiltonian into Eq. (S21) and after some massage, we can parameterize the pairing correlation as
+Substituting the BdG Hamiltonian into Eq. (S22) and after some algebraic manipulation, we can parameterize the pairing correlation as
 $F(\bm{k},i\omega_{n})=\Delta_{0}[C_{1}(\bm{k},i\omega_{n})+C_{2}(\bm{k},i\omega_{n})\bm{d}(\bm{k})\cdot\bm{s}]is_{y}$ (S23)
@@
 Therefore, it can be seen that due to the presence of $d$-wave, the spin-singlet pairing and the spin-triplet pairing are mixed.
-Microscopically, the spin-singlet pairing correlations provide the physical origin of the nonlinear spin magnetization.
-III III. Linear Edelstein effect in Rashba-type superconductors
-In this section we review the linear magnetoelectric effect in superconductors with Rashba SOC. The low energy Bloch Hamiltonian is
+Clearly, $C_{2}(\bm{k},i\omega_{n})=-C_{2}(-\bm{k},-i\omega_{n})$, which implies that the pairing amplitude changes sign under time-reversal operation.
+Therefore, the spin-triplet pairing is a type of odd-frequency pairing state Linder and Balatsky (2019).
+S3 Linear Edelstein effect in Rashba-type superconductors
+In this section, we review the linear magnetoelectric effect in superconductors with Rashba SOC. The low energy Bloch Hamiltonian is
 $H_{\bm{k}}=[ta^{2}(k_{x}^{2}+k_{y}^{2})-\mu]s_{0}+\lambda a(k_{x}s_{y}-k_{y}s_{x}),$ (S27)
@@
 Figure S5: The linear spin magnetization in Rashba-type superconductors.
-(a) The band structure with Rashba SOC at $\lambda=0.3t$.
+(a) The band structure with Rashba SOC at $\lambda=0.3t$ from Eq.
+(S27).
 (b) The first-order spin susceptibility $\alpha_{yx}$ as a function of $\lambda$ at $T=0.1T_{c}$, $\Delta_{0}=0.03t$ and $\mu=0.2t$.
@@
 (d) $\alpha_{yx}$ as a function of $T$ at $\Delta_{0}=0.1\lambda$ for various values of $\mu$.
-We apply the above formula to calculate the linear (first order) magnetization in superconductors with Rashba SOI.
+We apply the above formula to compute the linear (first order) magnetization in superconductors with Rashba SOI.
 The calculated band structure from Eq.
 (S27) is shown in Fig.
-S1(a) at $\lambda=0.3t$.
+S5(a) at $\lambda=0.3t$.
 By employing Eq.
-(S28), we plot the $\alpha_{yx}$ as a function of $\lambda$ in Fig.
-S1 (orange solid curve).
-By way of comparison, we also show a fully numerical plot of Eq.
+(S29), we plot $\alpha_{yx}$ as a function of $\lambda$ in Fig.
+S5(b) (orange solid curve).
+By way of comparison, we also show a numerical plot of Eq.
 (S11) (blue diamonds); both agree with each other.
 In Fig.
-S1(c) we show the $\mu$-dependence of the $\alpha_{yx}$ with various values of $\Delta_{0}$.
+S5(c) we show the $\mu$-dependence of $\alpha_{yx}$ with various values of $\Delta_{0}$.
 When $\mu<0$, $\alpha_{yx}$ decreases as $\mu$ goes down.
 Fig.
-S1(d) shows $\alpha_{yx}$ as a function of temperature for different values of $\mu$.
+S5(d) shows $\alpha_{yx}$ as a function of temperature for different values of $\mu$.
 Generally, $\alpha_{yx}$ starts to grow linearly when $T<T_{c}$, and it saturates when $T\rightarrow 0$.
-These behaviors are consistent with the results shown in Ref.
+This behavior is consistent with the results shown in Ref.
 He et al.
 (2019).
+S4 Symmetry analysis of NSM effect
+Supplementary Table S1: Symmetry restrictions of the NSM tensor $\chi_{ab}^{c}$ in a two-dimensional system.
+We have symmetrized $\chi_{xy}^{c}$ by defining $\chi_{(xy)}^{c}\equiv(\chi_{xy}^{c}+\chi_{yx}^{c})/2$.
+Here $\checkmark$ indicates allowed and $\times$ indicates forbidden.
+Here the /// symbol denotes “or”.
+$P$ $C_{2x}/\mathcal{M}_{x}$ $C_{2x}\mathcal{T}/\mathcal{M}_{x}\mathcal{T}$ $C_{2y}/\mathcal{M}_{y}$ $C_{2y}\mathcal{T}/\mathcal{M}_{y}\mathcal{T}$ $C_{2z}/\mathcal{M}_{z}$ $C_{2z}\mathcal{T}/\mathcal{M}_{z}\mathcal{T}$ $C_{3z}$ $C_{4z}/C_{6z}$ $C_{4z}\mathcal{T}$ $C_{6z}\mathcal{T}$ $C_{8z}\mathcal{T}/C_{12z}\mathcal{T}$
+$\chi_{xx}^{x}$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+$\chi_{(xy)}^{x}$$\times$ $\times$ $\times$ $-\chi_{yy}^{y}$$\times$ $\times$ $-\chi_{yy}^{y}$$\times$
+$\chi_{yy}^{x}$$\times$ $\times$ $\times$ $-\chi_{xx}^{x}$$\times$ $\times$ $-\chi_{xx}^{x}$$\times$
+$\chi_{xx}^{y}$$\times$ $\times$ $\times$ $-\chi_{yy}^{y}$$\times$ $\times$ $-\chi_{yy}^{y}$$\times$
+$\chi_{(xy)}^{y}$$\times$ $\times$ $\times$ $-\chi_{xx}^{x}$$\times$ $\times$ $-\chi_{xx}^{x}$$\times$
+$\chi_{yy}^{y}$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+$\chi_{xx}^{z}$$\times$ $\times$ $\times$ $\times$ $\times$
+$\chi_{(xy)}^{z}$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+$\chi_{yy}^{z}$$\times$ $\times$ $\times$ $\chi_{xx}^{z}$ $\chi_{xx}^{z}$ $-\chi_{xx}^{z}$$\times$ $\times$
+S4.1 Symmetry analysis
+In this section, we study the symmetry requirements for NSM (i.e.
+obtaining a non-zero $\chi_{ab}^{c}$) focussing on systems with time-reversal $\mathcal{T}$ breaking.
+Under a (composite) operation $\mathcal{T}^{s_{\mathcal{T}}}\mathcal{R}$, the NSM tensor $\chi_{ab}^{c}$ transforms as a third-rank pseudotensor, which can be written as
+$\chi_{a^{\prime}b^{\prime}}^{c^{\prime}}=(-1)^{s_{\mathcal{T}}}\mathrm{det}(\mathcal{R})\mathcal{R}_{c^{\prime}c}\mathcal{R}_{b^{\prime}b}\mathcal{R}_{a^{\prime}a}\chi_{ab}^{c}$ (S31)
+where $\mathcal{R}$ is the point group operation.
+Here we have used a short form notation to indicate composite operation.
+For point group $\mathcal{R}$ operation only, $s_{\mathcal{T}}=0$.
+For composite $\mathcal{R}\mathcal{T}$ operation, $s_{\mathcal{T}}=1$.
+We examine the symmetry requirements of NSM, listing how various (composite) point group symmetries affect the presence/absence of NSM as shown in Table.
+S1.
+From this symmetry analysis, we can clearly see that NSM can arise for a range of symmetries beyond $C_{4}\mathcal{T}$ that we focus on in the main text.
+It is also worthwhile to examine the presence/absence of NSM response for other types of altermagnetic order.
+For example, $g$-wave altermagnetic order $J_{ex}k_{x}k_{y}(k_{x}^{2}-k_{y}^{2})s_{z}$ and $i$-wave altermagnetic order $J_{ex}k_{x}k_{y}(3k_{x}^{2}-k_{y}^{2})(3k_{y}^{2}-k_{x}^{2})s_{z}$mejkal et al.
+(2022c).
+It is clear that for altermagnetic $g$-wave with $C_{8z}\mathcal{T}$ symmetry or $i$-wave with $C_{12z}\mathcal{T}$ symmetry, $\chi_{ab}^{c}=0$.
+S4.2 A case study: NSM in kagome noncollinear antiferromagnet
+In the main text, we focussed on the nonlinear superconducting magnetoelectric (NSM) effect in $d$-wave altermagnet/superconductor heterostructure.
+The formula we derive in Eq.
+(S12) is general and can also be applied to other magnet/superconductor heterostructure or intrinsic superconductor with magnetism.
+Indeed, in the symmetry analysis of the previous section, we find that NSM can appear for a range of other symmetries.
+In this section we present more details of another example (briefly described in the main text): Noncollinear antiferromagnet with $D_{3d}$ symmetry on kagome lattice Lee et al.
+(2024a).
+Figure S6: (a) The crystal structure for the kagome lattice.
+The noncollinear magnetic order is labeled by red arrows.
+(b) The energy spectrum of kagome lattice from Eq.
+(S32).
+(c) The BdG spectrum of the kagome oncollinear antiferromagnet [Eq.
+(S36)] with the pairing potential $\Delta_{0}is_{y}$.
+(d) The spin magnetization $M_{y}$ as a function of $q_{x}$ from full numerical method in Eq.
+(S6) (solid orange).
+As a comparison, the NSM susceptibility multiplied by $q_{x}^{2}$ is plotted as Eq.
+(S12) (dashed green).
+Parameters in the calculation:$t=1,m=0.2,\mu=0.1,\Delta_{0}=0.05$ and $T=0.3T_{c}$.
+Considering a kagome lattice with $A,B,C$ atoms per uint cell, in the basis $c_{\bm{k}}^{\dagger}=(c_{A\bm{k}}^{\dagger},c_{B\bm{k}}^{\dagger},c_{C\bm{k}}^{\dagger})$ the tight binding Hamiltonian is given by
+$H_{kagome}(\bm{k})=\begin{pmatrix}0&h_{12}&h_{13}\\ h_{12}^{*}&0&h_{23}\\ h_{13}^{*}&h_{23}^{*}&0\end{pmatrix}$ (S32)
+where $h_{12}=-(1+e^{i\bm{k}\cdot\bm{a_{1}}})$, $h_{13}=-(1+e^{-i\bm{k}\cdot\bm{a_{3}}})$, $h_{23}=-(1+e^{i\bm{k}\cdot\bm{a_{2}}})$ with $a_{1}=(1,0)a$, $a_{2}=(-1/2,\sqrt{3}/2)a$, $a_{3}=(-1/2,-\sqrt{3}/2)a$.
+The energy spectrum is plotted in Fig.
+S6(b).
+Next we consider the noncollinear magnetic order $h_{afm}=M_{s}(M_{x}s_{x}+M_{y}s_{y})$ [Fig.
+S6(a)], with
+$M_{x}=\begin{pmatrix}-\frac{\sqrt{3}}{2}&0&0\\ 0&\frac{\sqrt{3}}{2}&0\\ 0&0&0\end{pmatrix},M_{y}=\begin{pmatrix}-\frac{1}{2}&0&0\\ 0&-\frac{1}{2}&0\\ 0&0&1\end{pmatrix}.$ (S33)
+Here $M_{s}$ denotes the strength of noncollinear magnetic order.
+Considering the spin degree of freedom, the model Hamiltonian has six band.
+Here we consider the low-energy state near $\Gamma$ pocket in Fig.
+S6(b).
+To further facilitate analysis, we can make the standard perturbation theory analysis focussing on the wavefunction of the lowest band near $\Gamma$ pocket
+$|\psi\rangle\approx[1+\frac{i\sqrt{3}}{4}k_{y}-\frac{k_{x}^{2}+k_{y}^{2}}{16}+(\frac{i}{4}-\frac{\sqrt{3}}{12}k_{y})k_{x},1+\frac{i\sqrt{3}}{4}k_{y}-\frac{k_{x}^{2}+k_{y}^{2}}{16}+(-\frac{i}{4}+\frac{\sqrt{3}}{12}k_{y})k_{x},1]^{T}.$ (S34)
+Then we can obtain the momentum-space $h_{afm}(\bm{k})$ as
+$h_{afm}(\bm{k})=\langle\psi|h_{afm}|\psi\rangle/\langle\psi|\psi\rangle\approx\frac{1}{48}M_{s}[2k_{x}k_{y}s_{x}+(k_{x}^{2}-k_{y}^{2})s_{y}]$ (S35)
+One can find that the noncollinear magnetic order in real space gives rise to the momentum dependent magnetic order.
+Similar results are also obtained in Ref.
+Lee et al.
+(2024a).
+To study the NSM in the superconducting state, we adopt the low-energy effective Hamiltonian as
+$H_{\bm{k}}=[ta^{2}(k_{x}^{2}+k_{y}^{2})-\mu]s_{0}+2ma^{2}k_{x}k_{y}s_{x}+ma^{2}(k_{x}^{2}-k_{y}^{2})s_{y},$ (S36)
+where $t$ is the hopping term, $m$ denotes the noncollinear magnetic order, and $s_{x}$, $s_{y}$ are spin Pauli matrices.
+The Hamiltonian possesses $D_{3d}$ symmetry, with $C_{3z}=(s_{0}-i\sqrt{3}s_{z})/2$, $C_{2y}=is_{y}$ and $I=s_{0}$ (inversion) symmetries.
+It is straightforward to obtain the eigenenergy as $E_{\pm}=ta^{2}\bm{k}^{2}-\mu\pm ma^{2}\bm{k}^{2}$ and the corresponding spin expectation value as
+$(\langle s_{x}\rangle,\langle s_{y}\rangle)=(\pm\sin 2\theta,\pm\cos 2\theta),$ (S37)
+where $\theta$ is the polar angle.
+Based on the symmetry analysis in Table.
+S1, the nonzero spin susceptibilities are $\chi_{xx}^{y}=\chi_{xy}^{x}=\chi_{yx}^{x}=-\chi_{yy}^{y}$.
+Without loss of generality, we also adopt the spin-singlet pairing potential as $\hat{\Delta}=\Delta_{0}is_{y}$.
+The superconducting state can be intrinsic or proximity induced.
+In our calculations, we also adopt the BCS temperature dependence of $\Delta_{0}$ with $\Delta(T)=\Delta_{0}\tanh(1.74\sqrt{T_{c}/T-1})$ and $\Delta_{0}=1.76k_{B}T_{c}$.
+The results are shown in Fig.
+S6 with the model parameters $t=1,m=0.2,\mu=0.1$ and $\Delta_{0}=0.05$ and $T=0.3T_{c}$.
+In Fig.
+S6(c) we plot the BdG spectrum.
+Importantly, we plot Fig.
+S6(d) that displays $M_{y}$ from a numerical evaluation of Eq.
+(S6) as well as $\chi_{xx}^{y}q_{x}^{2}$ evaluated from the NSM susceptibility tensor $\chi_{ab}^{c}$ [Eq.
+(S12)].
+These show that an in-plane NSM manifests for this model.